### v1

# Chest X-ray Foundation Model with Global and Local Representations Integration

Abstract
Chest X-ray (CXR) is the most frequently ordered imaging test, supporting diverse clinical tasks from thoracic disease detection to postoperative monitoring. However, task-specific classification models are limited in scope, require costly labeled data, and lack generalizability to out-of-distribution datasets. To address these challenges, we introduce CheXFound, a self-supervised vision foundation model that learns robust CXR representations and generalizes effectively across a wide range of downstream tasks. We pretrain CheXFound on a curated CXR-1M dataset, comprising over one million unique CXRs from publicly available sources. We propose a Global and Local Representations Integration (GLoRI) module for downstream adaptations, by incorporating disease-specific local features with global image features for enhanced performance in multilabel classification. Our experimental results show that CheXFound outperforms state-of-the-art models in classifying 40 disease findings across different prevalence levels on the CXR-LT 24 dataset and exhibits superior label efficiency on downstream tasks with limited training data. Additionally, CheXFound achieved significant improvements on new tasks with out-of-distribution datasets, including opportunistic cardiovascular disease risk estimation and mortality prediction. These results highlight CheXFound’s strong generalization capabilities, enabling diverse adaptations with improved label efficiency. The project source code is publicly available at https://github.com/RPIDIAL/CheXFound.
{IEEEkeywords}
Chest X-ray, Foundation Model, Knowledge Distillation, Self-supervised Learning, Pretraining.
1 Introduction
[…]
Chest X-ray (CXR) is one of the most commonly-ordered imaging tests worldwide [1]. Clinical CXR interpretation encompasses a broad spectrum of tasks, including detecting diseases associated with the lungs, heart, blood vessel, and bones, as well as monitoring postoperative recovery and the positioning of support devices. With advancements in computer-aided diagnosis, these tasks now extend even further to include opportunistic disease risk assessment, such as cardiovascular disease[2, 3], mortality risk[4], and diabetes[5], among other factors not directly quantifiable by human eyes. Training specialized classification models for each task from scratch poses significant limitations. Such models are typically effective only within a narrow scope of pathologies and struggle to generalize to out-of-distribution datasets. Furthermore, developing these models requires requires extensive labeled datasets, which are both cost-prohibitive and inefficient. These challenges underscore the need for self-supervised models that can learn robust representations and demonstrate superior generalization capabilities across diverse tasks.
Recent advancements in the field of computer vision[6, 7, 8, 9] demonstrate that self-supervised vision models can produce task-agnostic and semantic-rich image representations that achieve improved performance on a broad spectrum of downstream tasks. Such models are called foundation models because of their superior capabilities to adapt to diverse downstream tasks when pretrained on large-scale data. Recent works in self-supervised learning for CXR interpretation adopt a series of advanced training strategies to learn high-quality image representations, including contrastive learning [10], masked image modeling (MIM) [11], and self-distillation [12]. Research further use CXRs and their clinical reports to perform contrastive language-image pretraining [13]. However, these studies have two major limitations. First, they only evaluate model performance for classifying a narrow range of disease findings, without considering the long-tail nature of pathologies in CXR and the opportunistic CXR interpretation tasks, such as cardiovascular disease (CVD) risk estimation and mortality prediction. Second, these studies simply rely on the global image features for disease classification, overlooking the use of CXR representations learned by the foundation models to provide disease-specific local features to enhance performance. Addressing these limitations is pivotal to the development of the CXR foundation models towards clinical applications which often involve interpreting a wide range of disease findings. It also has broader implications by enabling CVD risk estimation and mortality prediction with a routine CXR.
In this work, we introduce CheXFound, a vision foundation model specialized for CXR image analysis that learns high-quality CXR representations and generalizes effectively across a wide range of thoracic disease classification and opportunistic risk estimation tasks. We pretrain CheXFound on a curated CXR-1M dataset, comprising more than one million unique CXRs from 13 publicly available datasets, including MIMIC-CXR [14], CheXpert [15], PadChest [16], CXR14 [17], BRAX [18], VinDr-CXR [19], and CANDID-PTX [20], among others. Our CheXFound model is pretrained via DINOv2 [21], a state-of-the-art self-supervised learning method with strong off-the-shelf linear probe performance. For downstream adaptation, we propose a Global and Local Representations Integration (GLoRI) module. GLoRI is trained on top of the frozen CheXFound model. It uses the attentional principle to compute disease-specific local features and integrates them with the global image features to improve the multilabel classification performance.
We assess CheXFound’s performance on two tiers of CXR interpretation tasks, including thoracic disease classification and opportunistic risk estimation. CheXFound outperforms previous state-of-the-art models such as RAD-DINO [12], EVA-X [11], and CheXzero [13] across 40 disease findings at different prevalence levels on the CXR-LT 24 dataset [22]. Besides, CheXFound demonstrates superior label efficiency, which achieves best-performing results on the Shenzhen, Montgomery, and JSRT datasets with limited training data. We also find that CheXFound achieves significant performance increases compared with its comparisons for the out-of-distribution tasks, including opportunistic CVD risk estimation and mortality prediction on the PLCO dataset [23]. Overall, we demonstrate CheXFound’s strong generalization capabilities across a wide range of downstream tasks on in-distribution and out-of-distribution datasets. CheXFound’s strong representation quality can enable diverse downstream adaptations with improved label efficiency.
2 Related works
2.1 Self-supervised Visual Representation Learning
Our study is mostly related to self-supervised visual representation learning. After the success of masked language modeling in language domain, masked autoencoder (MAE) [7] and BEiT [24] translate the idea into visual representation learning, which assume the pretext task of recovering masked pixels can train networks to learn visual information and context. Another family of self-supervised learning methods (SimCLR [25] and MoCov3[6]) apply contrastive learning objectives, assuming augmentation invariance of image representations and aiming to learn contrastive class representations. These methods have been reported to achieve inferior linear probe performance and require fine-tuning backbone features [21]. They also do not translate well into medical applications [26]. Beyond the above methods, another family of self-supervised learning methods rely on a knowledge distillation framework first introduced by BYOL [27], which bootstraps latent features of a teacher network to train a student network. DINO [8] applies self-distillation with the Transformer architecture and enforce similarity of categorical distributions. iBOT [9] extends the framework with masked image modeling. DINOv2 [21] carefully curates pretraining data with deduplication and further makes modifications to improve training. Overall, self-distillation methods excel at linear probe evaluation and have demonstrated generalizability in medical application [12, 26]. Our study follows this methodology to train CheXFound with strong representation quality.
2.2 Foundation Models for Medical Applications
The surge in available data and computational resources have enabled the large-scale pretraining of foundation models. Studies have demonstrated that scaling foundation models in data and model sizes can achieve performance increases across a wide array of downstream tasks [7, 21, 26]. In medical domain, research works have developed multiple categories of foundation models differing in technical approaches and data modalities. Our study is related to vision-centric foundation models. RAD-DINO [12] and EVA-X [11] are two foundation models in CXR domain. Compared to CheXFound with ViT-L pretrained on CXR-1M, these models are limited in model and data scales. Another category of foundation models incorporate vision and text data for multimodal pretraining. CheXzero [13], BiomedCLIP [28], PubMedCLIP [29] use contrastive vision-language pretraining, which is effective in zero-shot classification. Further development of vision-language models takes advantage of instruction-tuning to improve reasoning and detailed description capabilities [30, 31, 32]. Overall, research empirically finds that vision-language models achieve inferior performance than vision-centric foundation models in CXR classification [12]. In this study, we focus on the vision-centric foundation model and investigate its capability for extensitve CXR classification tasks. To the best of our knowledge, our work employs the largest-scale self-supervised pretraining with over 1 million unique CXRs.
3 Materials and Methods
Figure 1: Overview of self-supervised pretraining of CheXFound, using publicly available CXRs from multiple institutions with a masked image modeling objective and a [CLS] token alignment objective.
Table 1: Curation of CXR-1M with publicly available datasets from diverse institutions for self-supervised pretraining. CXR-1M is subsetted into CXR-207K and CXR-744K to evaluate the data scalability of self-supervised models.
Datasets View Findings image #
MIMIC-CXR[14] Frontal, Lateral 14 diseases 207,096
Total number of images in CXR-207K: 207,096
CheXpert[15] Frontal, Lateral 14 diseases 223,648
PadChest[16] Frontal, Lateral 193 diseases 160,861
CXR14[17] Frontal 14 diseases 112,120
BRAX[18] Frontal, Lateal 14 diseases 40,967
Total number of images in CXR-744K: 744,692
VinDr-CXR[19] Frontal 28 diseases 18,000
CANDID-PTX[20] Frontal Pneumothorax 19,237
SIIM-ACR[33] Frontal Pneumothorax 18,499
Object-CXR[34] Frontal Foreign objects 9,000
COVID-19[35] Frontal COVID-19 7,597
COVIDx CXR-4[36] Frontal COVID-19 84,818
MIDRC COVIDx[37] Frontal COVID-19 23,001
BIMCV COVID+[38] Frontal, Lateral COVID-19 80,889
Total number of images in CXR-1M: 1,005,733
3.1 CXR-1M for Pretraining CheXFound
As detailed in Table 1, we curated the CXR-1M dataset for self-supervised pretraining by retrieving in total 1,005,733 unique CXRs from 13 publicly available datasets [14, 15, 16, 17, 36, 18, 37, 38, 19, 35, 34, 20, 33] that were released for various downstream tasks, including disease diagnosis, abnormality detection, foreign objection detection, and segmentation. To learn comprehensive representations for multiview CXR analysis, both frontal-view in PA (posterior-anterior) or AP (anterior-posterior) and lateral-view CXRs were included into CXR-1M.
To evaluate the data scalability of self-supervised pretraining, we further created CXR-207K and CXR-744K, denoting two subsets of CXR-1M, as shown in Table 1. CXR-207K contains approximately 207K CXRs from MIMIC-CXR. CXR-744K contains around 744K CXRs from five datasets: MIMIC-CXR, CheXpert, PadChest, CXR14, and BRAX.
We used DINOv2 [21], a state-of-the-art self-supervised learning method, to pretrain CheXFound on CXR-1M. DINOv2 inherits designs from DINO [8] and iBOT [9] and incorporates two self-distillation objectives: the masked image modeling loss $\mathcal{L}_{\texttt{MIM}}$ and the [CLS] token alignment loss $\mathcal{L}_{\texttt{[CLS]}}$. It uses a teacher-student knowledge distillation architecture as shown in Fig. 1 to learn CXR representations. Masked image modeling uses the teacher network as an online tokenizer, which generates patch tokens from intact images to guide the student network in reconstructing masked patch tokens. This approach enables the student network to learn both visual features and contextual information effectively. On the other hand, the [CLS] token alignment loss $\mathcal{L}_{\texttt{[CLS]}}$ enforces similarity between [CLS] tokens output by the teacher and student networks. This approach aims to train the network to learn high-level class representations with off-the-shelf linear probe capabilities.
3.2 Global and Local Representation Integration for Multilabel Classification
For the downstream evaluation of CheXFound, the linear probe classifier is a pivotal tool to evaluate the quality of pretrained representations. However, the linear probe classifier has limited capability to address the multilabel classification problem commonly seen in CXR interpretation, since it generally relies on the global image features from a single [CLS] token for classifying a wide range of pathologies and lacks essential local details to support the predictions. In contrast, patch tokens from our pretrained CheXFound contain rich CXR representations and high-level contextual information learned via masked image modeling, which can provide disease-specific local features to substantially reduce ambiguities arisen from using the [CLS] token for classifying multiple pathologies. To take advantage of both local and global features for disease classification, we introduce GLoRI (Fig. 2), which utilizes a cross-attention layer with disease queries to summarize patch token features and a skip-connection to integrate the [CLS] token towards final prediction.
To be specific, GLoRI receives output patch tokens $\mathbf{u}^{\texttt{Patch}}\in\mathbb{R}^{N\times D_{\text{model}}}$ from the frozen CheXFound backbone as input, where $D_{\text{model}}$ is the backbone embedding dimension. Since there can be a dimension mismatch between the backbone and GLoRI, we use a linear embedding layer to project $\mathbf{u}^{\texttt{Patch}}$ to the GLoRI dimensional space:
$\mathbf{u}^{\prime\texttt{Patch}}=\text{ReLU}(\text{Linear}^{\text{embed}}(\mathbf{u}^{\texttt{Patch}})),$ (1)
where $\mathbf{u}^{\prime\texttt{Patch}}\in\mathbb{R}^{N\times D_{\text{GLoRI}}}$ is the projected by the linear embedding layer $\text{Linear}^{\text{embed}}(\cdot)$ to the patch token sequence with dimension $D_{\text{GLoRI}}$. In GLoRI, to extract disease-specific local features using the cross-attention layer, we initialize $M$ disease queries corresponding to $M$ disease findings, denoted as $\mathbf{q}\in\mathbb{R}^{M\times D_{\text{key}}}$. The keys $\mathbf{k}\in\mathbb{R}^{M\times D_{\text{key}}}$ and values $\mathbf{q}\in\mathbb{R}^{M\times D_{\text{value}}}$ for the cross-attention layer are from $\mathbf{u}^{\prime\texttt{Patch}}$, which provides rich CXR representations. A scaled dot-product attention module is used to compute attention-pooled features relevant to the query diseases:
$\mathbf{q}^{\prime}=\text{Softmax}\left(\frac{\text{Linear}^{\text{query}}(\mathbf{q})\text{Linear}^{\text{key}}(\mathbf{k})^{T}}{\sqrt{D_{\text{key}}}}\right)\text{Linear}^{\text{value}}(\mathbf{v}),$ (2)
where $\mathbf{q}^{\prime}\in\mathbb{R}^{M\times D_{\text{key}}}$ denotes the output disease queries, which we consider contain disease-specific local features. $D_{\text{key}}$ is the dimension of the key tokens. Last, we concatenate $\mathbf{q}^{\prime}$ with the [CLS] token to construct the GLoRI output token sequence. For multilabel classification, each GLoRI output token is projected by a linear classifier supervised by a binary cross-entropy loss.
Overall, GLoRI applies the attentional principle, which shares similarity with previous works, such as Perceiver [39] and DETR [40], for CheXFound’s downstream evaluation. We empirically demonstrate that GLoRI extracts local features relevant to disease abnormalities in CXRs in Section 6.1.
Figure 2: Global and Local Representations Integration (GLoRI) for evaluating CheXFound on downstream tasks. GLoRI is appended on top of the frozen CheXFound backbone. GLoRI uses disease queries to compute attention-pooled CXR representations and integrates the global image features from [CLS] token for disease finding classification. For the opportunistic CXR interpretation, additional disease queries are created for GLoRI to retrieve representations related to the CVD risk estimation and mortality prediction.
4 Experimental Design
4.1 Implementation Details
4.1.1 Self-supervised pretraining details
We conducted self-supervised pretraining of CheXFound on our curated CXR-1M dataset (Table 1) using the ViT-L architecture with a patch size of 16$\times$16 pixels. The loss weights for $\mathcal{L}_{\texttt{[CLS]}}$ and $\mathcal{L}_{\texttt{MIM}}$ were set to 1.0 and 3.0, respectively. The momentum to compute the exponential moving average of the student network was set to 0.994. We varied the global and local crop sizes to pretrain CheXFound at a rang of resolutions. Specifically, we set the global and local crop size pairs to be (512, 144), (448, 128), (336, 128), and (224, 96). We set the number of global and local crops to 2 and 8, respectively. For masked image modeling, we set the proportion of masked patches to the range (0.1, 0.5). We trained CheXFound for 100 epochs with an epoch length of 2,500 iterations and a batch size of 14 per graphics processing unit (GPU). We used the AdamW optimizer with an initial learning rate of 2e-4. We applied a Cosine annealing schedule for learning rate decay and a warm up period of 10 epochs. We pretrained CheXFound on a DGX-1 server with 8×8$\times$8 × NVIDIA A100 40GB GPUs. Depending on the image resolutions, the pretraining processes take around 48 to 96 hours. The project source code is publicly available at https://github.com/RPIDIAL/CheXFound.
4.1.2 Downstream evaluation details
In downstream evaluation, we trained GLoRI with feature representations from the frozen CheXFound. We took the concatenated representations from the last 4 layers of CheXFound as the input to GLoRI. We set the embedding dimension of GLoRI to 768. Disease queries were randomly initialized with a standard normal distribution. For the cross-attention layer, we used the multihead attention mechanism and set the number of heads to 8. To train GLoRI, we used the AdamW optimizer and conduct a thorough learning rate search in {1e-5, 2e-5, 5e-5, 1e-4, 2e-4, 5e-4, 1e-3, 2e-3, 5e-3} to obtain the best-performing learning rate on the validation set. We then combined the training and validation sets for a second round of training using the best learning rate. To maximize the number of images that the GLoRI module processes during downstream adaptation, we trained GLoRI for 10 epochs on CXR-LT 24, CheXpert, and PLCO and 100 epochs on Shenzhen, Mongomery, and JSRT. We set the batch size to 16 for Montgomery and JSRT and 256 for CXR-LT 24, CheXpert, Shenzhen, and PLCO.
4.2 Experimental Design
To rigorously evaluate CheXFound’s in-distribution and out-of-distribution performance, we employed an extensive classification benchmark, consisting of in-distribution datasets (CXR-LT 24 [22] and CheXpert [15]) and out-of-distribution datasets (Shenzhen [41], Montgomery[41], Japanese Society of Radiological Technology (JSRT) [42], and Prostate, Lung, Colorectal, and Ovarian (PLCO) Cancer Screening Trial [23]). Following recent works in vision-centric foundation models [12, 26], we split the evaluation datasets into training, validation, and test splits. To avoid any potential data contamination, we included only the training set in CXR-LT 24 and the training and validation sets in CheXpert for self-supervised pretraining, while keeping the test set unseen. For the out-of-distribution datasets (Shenzhen, Montgomery, and JSRT), none of the images in the training, validation, and test sets were used for self-supervised pretraining.
Although the MIMIC-CXR dataset is the common benchmark to evaluate the performance of CXR interpretation models, its labels contain only 14 findings. To assess the generalizability of foundation models across diverse disease types, we conducted experiments on the CXR-LT 24 dataset which includes the annotations of 40 disease findings at different levels of prevalence (Fig. 4a). To evaluate the model performance, we divided CXR-LT 24 into a training set of 207,096 images and a test set of 51,775 images.
To evaluate the performance of the foundation models against the annotations on five selected pathologies (atelectasis, cardiomegaly, consolidation, edema, and pleural effusion) by board-certificated radiologists, we incorporated the CheXpert dataset. It was divided into 191,027 frontal-view images in the training set, 202 images in the validation set, and 518 images in the test set.
To assess the out-of-distribution generalization capabilities of the foundation models, we also performed evaluation on the Shenzhen and Montgomery datasets for tuberculosis detection and the JSRT database for lung nodule detection. The Shenzhen, Montgomery, and JSRT datasets were divided into training, validation, and test splits with a ratio of 70:10:20. Shenzhen contains 463 training images, 65 validation images, and 134 test images. Montgomery contains 96 training images, 14 validation images, and 28 test images. JSRT contains 171 training images, 24 validation images, and 50 test images.
To evaluate the extended predictive power of the foundation models, we obtained the lung screening CXRs from the PLCO trial and extracted the all-cause mortality and cardiovascular disease mortality labels from the up to 25-year follow-up data. The PLCO CXRs were divided into training, validation, and test sets of 133,543 images, 19,099 images, and 38,058 images respectively.
Since the CXR interpretation problem often involves severe class imbalance, we employed two metrics to evaluate model performance: the area under the precision-recall curve (AUPRC) and the area under the receiver operating characteristic curve (AUROC). For the multilabel classification problem, we computed the average of the metrics over disease findings. We estimated the 95% confidence intervals of the model performance in AUPRC and AUROC over 1,000 bootstrapped samples. To test statistical significance, we used a two-sided paired permutation test with 1,000 permutations to assess the observed performance differences of the two models for disease findings.
5 Experimental Results
Table 2: Comparison CheXFound and other foundation models when using linear probe and GLoRI for classification. Values inside the parentheses indicate the 95% confidence intervals. Values in bold indicate the best results.
Classifier Foundation CXR-LT 24 CheXpert Shenzhen Montgomery JSRT
methods models AUPRC AUROC AUPRC AUROC AUPRC AUROC AUPRC AUROC AUPRC AUROC
Linear probe PubMedCLIP [29] 0.089(.088–.089) 0.561(.554–.568) 0.277(.244–.311) 0.595(.564–.628) 0.857(.786–.913) 0.814(.738–.887) 0.565(.306–.817) 0.534(.310–.750) 0.785(.630–.912) 0.685(.511–.833)
BiomedCLIP [28] 0.117(.116–.118) 0.643(.636–.649) 0.557(.504–.607) 0.841(.822–.860) 0.903(.843–.949) 0.885(.827–.934) 0.925(.791–1.0) 0.929(.822–1.0) 0.589(.427–.755) 0.432(.265–.601)
CheXzero [13] 0.112(.111–.112) 0.552(.545–.558) 0.468(.424–.509) 0.778(.754–.798) 0.929(.884–.962) 0.906(.851–.950) 0.964(.869–1.0) 0.970(.898–1.0) 0.759(.597–.885) 0.620(.453–.777)
EVA-X [11] 0.114(.113–.115) 0.596(.590–.602) 0.468(.422–.512) 0.788(.763–.812) 0.840(.738–.921) 0.824(.746–.898) 0.577(.320–.809) 0.509(.278–.749) 0.641(.472–.797) 0.490(.312–.660)
RAD-DINO [12] 0.114(.113–.114) 0.557(.570–.583) 0.463(.422–.503) 0.746(.715–.778) 0.883(.818–.933) 0.861(.798–.916) 0.637(.369–.842) 0.561(.316–.788) 0.747(.585–.888) 0.623(.456–.778)
CheXFound 0.209(.204–.214) 0.799(.794–.804) 0.620(.630–.727) 0.876(.860–.892) 0.974(.949–.992) 0.967(.935–.990) 0.988(.939–1.0) 0.990(.952–1.0) 0.918(.826–.975) 0.856(.741–.948)
GLoRI head PubMedCLIP [29] 0.116(.115–.117) 0.649(.643–.655) 0.501(.450–.552) 0.804(.778–.828) 0.897(.832–.946) 0.867(.807–.927) 0.628(.363–.866) 0.694(.484–.872) 0.652(.483–.810) 0.493(.330–.654)
BiomedCLIP [28] 0.122(.121–.123) 0.643(.636–.649) 0.552(.506–.593) 0.829(.809–.847) 0.921(.870–.957) 0.897(.840–.944) 0.900(.738–1.0) 0.898(.744–1.0) 0.634(.470–.789) 0.505(.291–.704)
CheXzero [13] 0.131(.130–.132) 0.671(.665–.677) 0.599(.551–.647) 0.888(.868–.905) 0.912(.852–.954) 0.894(.838–.942) 0.726(.451–.901) 0.653(.396–.882) 0.640(.467–.807) 0.485(.323–.662)
EVA-X [11] 0.149(.147–.150) 0.679(.672–.685) 0.614(.571–.659) 0.870(.853–.888) 0.928(.881–.964) 0.896(.835–.945) 0.977(.909–1.0) 0.979(.918–1.0) 0.866(.760–.936) 0.748(.615–.864)
RAD-DINO [12] 0.173(.171–.176) 0.723(.717–.729) 0.639(.593–.687) 0.884(.869–.898) 0.909(.854–.952) 0.885(.823–.936) 0.909(.747–1.0) 0.911(.782–1.0) 0.683(.503–.861) 0.615(.440–.783)
CheXFound 0.252(.247–.258) 0.830(.826–.834) 0.679(.630–.727) 0.908(.894–.921) 0.983(.960–.996) 0.978(.951–.995) 1.000(1.0–1.0) 1.000(1.0–1.0) 0.986(.956–1.0) 0.975(.931–1.0)
Table 3: Comparison of CheXFound with the end-to-end trained model and vision-language foundation models. Values inside the parentheses indicate the 95% confidence intervals. Values in bold indicate the best results.
Classifier Foundation CXR-LT 24 CheXpert Shenzhen Montgomery JSRT
methods models AUPRC AUROC AUPRC AUROC AUPRC AUROC AUPRC AUROC AUPRC AUROC
End-to-end ConvNeXt [43] 0.170(.168–.173) 0.761(.756–.766) 0.657(.608–.704) 0.886(.867–.903) 0.923(.870–.963) 0.891(.823–.945) 0.622(.362–.862) 0.670(.458–.872) 0.750(.586–.886) 0.608(.456–.761)
Img-text align. PubMedCLIP [29] 0.068(.068–.069) 0.531(.524–.537) 0.255(.221–.294) 0.582(.550–.618) 0.577(.457–.700) 0.540(.437–.637) 0.440(.246–.672) 0.426(.216–.658) 0.596(.426–.758) 0.403(.244–.572)
BiomedCLIP [28] 0.071(.071–.072) 0.539(.533–.545) 0.356(.317–.393) 0.653(.626–.681) 0.795(.705–.878) 0.760(.680–.836) 0.786(.567–.936) 0.714(.487–.912) 0.692(.516–.859) 0.554(.397–.722)
CheXzero [13] 0.134(.133–.136) 0.668(.662–.674) 0.646(.600–.692) 0.888(.868–.905) 0.875(.804–.933) 0.849(.776–.911) 0.967(.870–1.0) 0.969(.889–1.0) 0.708(.532–.863) 0.530(.378–.698)
GLoRI head CheXFound 0.252(.247–.258) 0.830(.826–.834) 0.679(.630–.727) 0.908(.894–.921) 0.983(.960–.996) 0.978(.951–.995) 1.000(1.0–1.0) 1.000(1.0–1.0) 0.986(.956–1.0) 0.975(.931–1.0)
5.1 Overall performance comparison
A pivotal characteristic of foundation models lies in their capability to achieve improved performance on a wide range of downstream datasets. To evaluate the capability of foundation models, we compared CheXFound, which uses ViT-L pretrained on CXR-1M, with publicly available pretrained encoders, including RAD-DINO [12], EVA-X [11], CheXzero [13], BiomedCLIP [28], PubmedCLIP [29], and ConvNeXt [43]. RAD-DINO was pretrained on a combined dataset comprising MIMIC-CXR, CheXpert, PadChest, CXR14, and BRAX, using the DINOv2 framework. EVA-X was pretrained on MIMIC-CXR, CheXpert, and CXR14 using EVA [44] technique with contrastive vision features for masked image modeling. CheXzero, a vision-language foundation model, was pretrained on MIMIC-CXR using contrastive language-image pretraining (CLIP) [45]. BiomedCLIP was pretrained on PMC-15M with image-text pairs collected from scientific articles. PubmedCLIP was initialized with the CLIP model and finetuned on PubMed articles. Finally, ConvNeXt was pretraiend on the ImageNet-22K dataset.
Figure 3: Detailed Performance for 40 disease findings in AUROC on the CXR-LT 24 dataset. Our CheXFound is compared with the vision-centric foundation models (EVA-X and RAD-DINO), the vision-language pretrained foundation models (CheXzero, BiomedCLIP, and PubMedCLIP), and the end-to-end trained model (ConvNeXt) with ImageNet-22K pretraining.
Figure 4: Model performance under high, medium and low disease prevalence. a, The number of labels for the 40 disease findings on the CXR-LT 24 dataset [22]. b, Model performance in AUPRC stratified by high, medium and low disease prevalence. Error bars indicate the 95% confidence intervals of AUPRC over 1,000 bootstrapped samples.
To evaluate the effectiveness and generalizability of the representations extracted by the foundation models, we evaluated the linear probe performance across the five datasets (CXR-LT 24, CheXpert, Shenzhen, Montgomery, and JSRT). While linear probe provides a straightforward approach to evaluate the quality of representation, it primarily relies on global image features from the [CLS] token, often resulting in suboptimal performance. Hence, we also evaluated the pretrained encoders using GLoRI, which incorporates attention-pooled local features in addition to the global image features from the [CLS] token. For vision-language pretained encoders, we further validated the quality of their vision representations by examing their correlation with text features for disease finding classification.
Across all the five classification tasks (CXR-LT 24, CheXpert, Shenzhen, Montgomery, and JSRT), CheXFound consistently outperformed other foundation models in the linear probe setting as shown in Table 2. On the multilabel, long-tailed classification task (CXT-LT 24), CheXFound achieved an AUPRC of 0.209, outperforming the next best-performing model (either RAD-DINO or EVA-X) by 9.5% ($p<0.001$, two-sided paired permutation test). On the five-class multilabel classification task (CheXpert), CheXFound outperformed the next best-performing model (EVA-X) by 8.8% ($p<0.001$) in AUROC. On single-class classification tasks (Shenzhen, Montgomery, and JSRT) with limited amounts of training data, CheXFound similarly outperformed the next best-performing models in AUROC by 11.7% ($p<0.001$), 42.9% ($p<0.001$), and 23.3% ($p<0.001$), respectively.
We further evaluated the performance of the foundation models using GLoRI across five classification tasks (Table 2). The foundation models (CheXFound, RAD-DINO, EVA-X, CheXzero, BiomedCLIP, and PubMedCLIP) with GLoRI generally outperformed their linear probe baselines. Specifically, CheXFound with GLoRI outperformed its linear probe baseline in AUROC by 3.1%, 3.2%, 1.1%, 1.0%, and 11.9% on CXR-LT 24, CheXpert, Shenzhen, Montgomery, and JSRT, respectively. CheXFound with GLoRI also outperformed other foundation models, including CLIP-based models (CheXzero, BioMedCLIP, and PubMedCLIP) under image-text alignment, as well as end-to-end trained ConvNeXt (Table 3).
To show the detailed performance of CheXFound over 40 disease findings, we illustrated CheXFound performance in AUROC against its comparisons in Fig. 3. CheXFound consistently outperformed other methods across the 40 disease findings in AUROC. We also compared performance in AUPRC over disease findings with high, medium, and low prevalence in Fig. 3b. CheXFound outperformed its comparisons under all levels of prevalence, even for underrepresented pathologies in the low prevalence category.
5.2 Opportunistic predictive power
Beyond thoracic disease detection tasks, we evaluated ChesXFound’s generalizability in opportunistic CXR interpretation. For this purpose, we requested access to the CXR arm of the PLCO trial [46], which includes digitally scanned CXR films and up to 25-year morality follow-up data. Using this dataset, we investigated CheXFound’s predictive capability for cardiovascular disease (CVD) risk and all-cause mortality estimation. We used CheXFound with ViT-L pretrained on CXR-1M in this experiment and compared CheXFound against two vision foundation models (RAD-DINO, EVA-X) and the end-to-end trained model, ConvNeXt.
CheXFound consistently outperformed its counterparts in both CVD risk and all-cause mortality estimation tasks. CheXFound achieved 0.749 for CVD risk estimation and 0.786 for all-cause mortality estimation in AUROC, significantly outperforming the next best-performing method (ConvNeXt) by 3.5% ($p<0.001$) and 4.1% ($p<0.001$), respectively. For the all-cause mortality estimation task, we divided the test cohort into low-risk and high-risk groups based on the model prediction and computed their Kaplan-Meier curves (Fig. 5). The survival distributions for the low-risk and high-risk groups are statistically different ($p<0.001$, log-rank test). The end-point survival probabilities are also different by a large margin for low-risk and high-risk groups (78.4% versus 38.4%). Overall, we demonstrated CheXFound generalization capability for opportunistic CXR interpretation.
Table 4: Model performance on CVD risk and all-cause mortality estimation. Values inside the parentheses are 95% confidence intervals. Values in bold indicate the best-performing results.
Methods CVD risk All-cause mortality
AUPRC AUROC AUPRC AUROC
ConvNeXt [43] 0.249(.240–.257) 0.714(.705–.723) 0.638(.629–.646) 0.745(.736–.753)
EVA-X [11] 0.179(.171–.188) 0.643(.635–.652) 0.545(.536–.554) 0.680(.674–.686)
RAD-DINO [12] 0.223(.215–231) 0.687(.679–.695) 0.615(.607–.622) 0.723(.716–.729)
CheXFound 0.289(.276–.301) 0.749(.741–.756) 0.695(.687–.702) 0.786(.782–.791)
Figure 5: Kaplan-Meier curves for low-risk and high-risk groups of all-cause mortality on the PLCO dataset. The number of subjects in the test set is 10,509. The survival rates for low-risk and high-risk groups are significantly different (p$<$0.0001, log-rank test). Shaded areas indicate 95% confidence intervals.
Table 5: Model Performance across pretraining data sizes and model scales. Results are given in the mean values of AUPRC and AUROC over 1000 bootstrapped samples. Values inside the parentheses indicate the 95% confidence intervals. Values in bold indicate the best-performing results.
Pretrain. Arch. CXR-LT 24 CheXpert Shenzhen Montgomery JSRT
Data AUPRC AUROC AUPRC AUROC AUPRC AUROC AUPRC AUROC AUPRC AUROC
CXR-207K ViT-Base 0.185(.182–.188) 0.775(.753–.789) 0.598(.550–.645) 0.854(.839–.862) 0.921(.898–.946) 0.909(.855–.953) 0.889(.732–1.0) 0.897(.742–1.0) 0.756(.591–.882) 0.612(.495–.712)
ViT-Large 0.207(.203–.210) 0.795(.774–.813) 0.618(.629–.726) 0.874(.862–.884) 0.953(.923–.979) 0.940(.901–.962) 0.923(.811–1.0) 0.932(.826–1.0) 0.775(.620–.892) 0.668(.537–.784)
CXR-744K ViT-Base 0.211(.208–.214) 0.803(.792–.810) 0.614(.571–.659) 0.869(.854–.886) 0.938(.913–.961) 0.925(.875–.963) 0.909(.775–1.0) 0.915(.785–1.0) 0.866(.759–.935) 0.747(.695–.787)
ViT-Large 0.217(.214–.221) 0.813(.801–.822) 0.643(.596–.692) 0.887(.876–.896) 0.964(.935–.986)) 0.956(.923–.972) 0.953(.843–1.0) 0.957(.845–1.0) 0.905(.794–.953) 0.826(.674–.925)
CXR-1M ViT-Base 0.219(.215–.223) 0.815(.811–.819) 0.632(.583–.679) 0.877(.853–.896) 0.957(.924–.979) 0.947(.925–.963) 0.923(.788–1.0) 0.927(.818–1.0) 0.908(.806–.955) 0.845(.696–.943)
ViT-Large 0.252(.247–.258) 0.830(.826–.834) 0.679(.630–.727) 0.908(.894–.921) 0.983(.960–.996) 0.978(.951–.995) 1.000(1.0–1.0) 1.000(1.0–1.0) 0.986(.956–1.0) 0.975(.931–1.0)
5.3 Scalability of self-supervised vision encoders
The scaling capabilities of the self-supervised vision encoders depends on both the model size and the pretraining data size and diversity [21, 26]. To analyze the scaling trends, we pretrained CheXFound across a range of data scales, with CXR-1M and its two subsets CXR-744K and CXR-207K. We also evaluated the impact of model scale by using ViT-Base (ViT-B) and ViT-Large (ViT-L) as the backbones.
Our results in Table 5 demonstrate that CheXFound benefits form both data and model scaling. Increasing the pretraining data from CXR-207K to CXR-1M with a ViT-L backbone leads to significant AUROC improvements of 3.5% ($p<0.001$) on CXR-LT 24, 3.4% ($p<0.001$) on CheXpert, 3.8% ($p<0.001$) on Shenzhen, 6.8% ($p<0.001$) on Montgomery, and 30.7% ($p<0.001$) on JSRT. We observed similar trends when using ViT-B, with performance also improving as we scale up the pretraining data from CXR-207K to CXR-1M. In addition, CheXFound with ViT-L consistently outperformed the ViT-B architecture across different data sizes. These results align with previous studies on scaling ViT models [21, 12, 26].
Figure 6: Evaluation of CheXFound model on the CXR-LT 24, CheXpert, Shenzhen, Montgomery, and JSRT datasets across a range of pretraining image resolutions. Error bars indicate 95% confidence intervals.
Figure 7: GLoRI attention maps for disease findings of atelectasis, cardiomegaly, consolidation, edema, and pleural effusion, respectively. Each subfigure contains 25 CXRs with their GLoRI attention maps overlaid (bottom) and an anchor CXR with a global attention map overlaid (top). The global attention maps are the averaged attention maps of 25 CXRs after registered to the anchor image.
5.4 Impact of CXR resolution
To assess the impact of CXR resolution used for pretraining, we pretrained CheXFound using ViT-L with a patch size of 16 across a range of input resolution, including $224^{2}$, $336^{2}$, $448^{2}$, and $512^{2}$. We empirically found that pretraining at high resolution from scratch cannot produce meaningful representations for downstream tasks. To deal with this problem, we pretrained CheXFound at resolution $224^{2}$ from scratch and then used the pretrained weights to initialize higher-resolution pretraining at $336^{2}$, $448^{2}$, and $512^{2}$.
Fig. 6 shows that self-supervised pretraining at higher resolutions results in improved performance on downstream tasks. Increasing the pretraining resolutions from 2242 pixels to 5122 pixels significantly improves the AUROC by 3.3% ($p<0.001$), 3.5% ($p<0.001$), 3.7% ($p<0.001$), 10.7% ($p<0.001$), and 35.8% ($p<0.001$) on the CXR-LT 24, CheXpert, Shenzhen, Montgomery, and JSRT datasets, respectively.
6 Discussion and Conclusion
6.1 Interpretation of disease-specific local features
The interpretabliliy of an artificial intelligence model is crucial to its medical applications. In this study, we trained GLoRI on top of the frozen foundation model. GLoRI inherently provides interpretable attention maps for each pathology. We visualized the attention maps for a selection of five pathologies of 25 CXRs and aligned these attention maps to an anchor CXR via affine registration [^1] We apply affine registration using the SimpleElastix library: https://simpleelastix.readthedocs.io. to provide a global perspective in Fig. 7. The attention maps contain precise localization of abnormalities in CXRs and the global attention maps cover the regions where the pathologies constantly occur. To be specific, edema refers to the accumulation of excess fluid and its abnormal regions often diffuse across lungs. This pattern is well captured by our attention maps as we observed in Fig. 7d that the critical regions in the individual attention maps scatter over the lungs and the global attention map covers extensive regions of both lungs. However, these attention maps have the limitations of covering partial regions of the abnormalities. For example, the attention maps for cardiomegaly only cover the heart on the left and right regions of the spine, and some maps for pleural effusion only cover the inferior boundaries of the lung while ignore the remaining abnormal regions.
6.2 Generalizability of foundation models
An important characteristic of CheXFound and other foundation models is their generalization capabilities to in-distribution and out-of-distribution downstream tasks. Compared with other encoders, we found that CheXFound achieved better performances on both in-distribution CXR-LT 24 and CheXpert datasets and out-of-distribution Shenzhen, Montgomery, and JSRT datasets. On the opportunistic CXR interpretation tasks on PLCO, CheXFound also achieved consistent and significant increases over comparison methods. CheXFound’s generalizability is attributed to the strong representation quality of frozen features learned via pretraining with large-scale, diverse CXRs. We also demonstrated CheXFound’s generalization capabilities on infrequent and underrepresented pathologies. CheXFound achieved significant increases over its comparions in classifying low-prevalence pathologies with lower than 1% occurrence frequencies, demonstrating its superior label efficiency. Although CheXFound with ViT-L achieved robust generalizability, our study did not evaluate the best-performing ViT-giant (ViT-g) architecture in DINOv2, a larger model with 1.1B parameters, which we expect to achieve better generalization performances in CXR interpretation, but it demands more pretraining data and computational resources. Overall, we demonstrated CheXFound’s robust generalization capabilities, which we believe can enable diverse downstream adaptations with improved label efficiency.
6.3 Conclusion
In summary, this work introduces CheXFound, a vision-centric foundation model pretrained via self-distillation on over one million unique CXRs. For downstream adaptations, we trained a GLoRI module on top of the frozen CheXFound, which combines disease-specific local features and global image features to improve the multilabel classification performance. CheXFound outperformed previous methods for classifying 40 disease findings on CXR-LT 24, demonstrating superior label efficiency on datasets with limited training labels, and strong generalization capabilities for opportunistic CXR interpretation on PLCO. The disease-specific local features extracted from CheXFound also carry strong interpretability as visuliazed by the associated attention maps. In our future work, we will continue to explore novel pretraining schemes to further improve the understanding of CXRs by these foundation models.

### v2

@@
 To address these challenges, we introduce CheXFound, a self-supervised vision foundation model that learns robust CXR representations and generalizes effectively across a wide range of downstream tasks.
-We pretrain CheXFound on a curated CXR-1M dataset, comprising over one million unique CXRs from publicly available sources.
-We propose a Global and Local Representations Integration (GLoRI) module for downstream adaptations, by incorporating disease-specific local features with global image features for enhanced performance in multilabel classification.
-Our experimental results show that CheXFound outperforms state-of-the-art models in classifying 40 disease findings across different prevalence levels on the CXR-LT 24 dataset and exhibits superior label efficiency on downstream tasks with limited training data.
-Additionally, CheXFound achieved significant improvements on new tasks with out-of-distribution datasets, including opportunistic cardiovascular disease risk estimation and mortality prediction.
-These results highlight CheXFound’s strong generalization capabilities, enabling diverse adaptations with improved label efficiency.
+We pretrained CheXFound on a curated CXR-987K dataset, comprising over approximately 987K unique CXRs from 12 publicly available sources.
+We propose a Global and Local Representations Integration (GLoRI) head for downstream adaptations, by incorporating fine- and coarse-grained disease-specific local features with global image features for enhanced performance in multilabel classification.
+Our experimental results showed that CheXFound outperformed state-of-the-art models in classifying 40 disease findings across different prevalence levels on the CXR-LT 24 dataset and exhibited superior label efficiency on downstream tasks with limited training data.
+Additionally, CheXFound achieved significant improvements on downstream tasks with out-of-distribution datasets, including opportunistic cardiovascular disease risk estimation, mortality prediction, malpositioned tube detection, and anatomical structure segmentation.
+The above results demonstrate CheXFound’s strong generalization capabilities, which will enable diverse downstream adaptations with improved label efficiency in future applications.
 The project source code is publicly available at https://github.com/RPIDIAL/CheXFound.
@@
 Such models are typically effective only within a narrow scope of pathologies and struggle to generalize to out-of-distribution datasets.
-Furthermore, developing these models requires requires extensive labeled datasets, which are both cost-prohibitive and inefficient.
+Furthermore, developing these models requires extensive labeled datasets, which are both cost-prohibitive and inefficient.
 These challenges underscore the need for self-supervised models that can learn robust representations and demonstrate superior generalization capabilities across diverse tasks.
@@
 However, these studies have two major limitations.
-First, they only evaluate model performance for classifying a narrow range of disease findings, without considering the long-tail nature of pathologies in CXR and the opportunistic CXR interpretation tasks, such as cardiovascular disease (CVD) risk estimation and mortality prediction.
-Second, these studies simply rely on the global image features for disease classification, overlooking the use of CXR representations learned by the foundation models to provide disease-specific local features to enhance performance.
+First, most of these proposed models are constrained by either model capacity or pretraining data size, limiting their capability to learn rich and effective representations for long-tail classification, as demonstrated by our detailed experimental results in Fig.
+3, showing that existing foundation models achieved degraded performance across all 40 diseases compared to CheXFound with large model capacity and data size.
+Second, most of these studies simply rely on the global image features for disease detection, overlooking the use of patch embeddings learned by the foundation model to provide disease-specific local features to enhance performance.
 Addressing these limitations is pivotal to the development of the CXR foundation models towards clinical applications which often involve interpreting a wide range of disease findings.
 It also has broader implications by enabling CVD risk estimation and mortality prediction with a routine CXR.
-In this work, we introduce CheXFound, a vision foundation model specialized for CXR image analysis that learns high-quality CXR representations and generalizes effectively across a wide range of thoracic disease classification and opportunistic risk estimation tasks.
-We pretrain CheXFound on a curated CXR-1M dataset, comprising more than one million unique CXRs from 13 publicly available datasets, including MIMIC-CXR [14], CheXpert [15], PadChest [16], CXR14 [17], BRAX [18], VinDr-CXR [19], and CANDID-PTX [20], among others.
-Our CheXFound model is pretrained via DINOv2 [21], a state-of-the-art self-supervised learning method with strong off-the-shelf linear probe performance.
-For downstream adaptation, we propose a Global and Local Representations Integration (GLoRI) module.
-GLoRI is trained on top of the frozen CheXFound model.
-It uses the attentional principle to compute disease-specific local features and integrates them with the global image features to improve the multilabel classification performance.
-We assess CheXFound’s performance on two tiers of CXR interpretation tasks, including thoracic disease classification and opportunistic risk estimation.
-CheXFound outperforms previous state-of-the-art models such as RAD-DINO [12], EVA-X [11], and CheXzero [13] across 40 disease findings at different prevalence levels on the CXR-LT 24 dataset [22].
-Besides, CheXFound demonstrates superior label efficiency, which achieves best-performing results on the Shenzhen, Montgomery, and JSRT datasets with limited training data.
-We also find that CheXFound achieves significant performance increases compared with its comparisons for the out-of-distribution tasks, including opportunistic CVD risk estimation and mortality prediction on the PLCO dataset [23].
+In this work, we introduce CheXFound, a vision foundation model specialized for CXR image analysis that learns high-quality CXR representations and generalizes effectively across a wide range of thoracic disease detection and opportunistic risk estimation tasks.
+We pretrained CheXFound on a curated CXR-987K dataset, comprising approximately 987K unique CXRs from 12 publicly available datasets, including MIMIC-CXR [14], CheXpert [15], PadChest [16], NIH-CXR [17], BRAX [18], and CANDID-PTX [19], among others.
+Our CheXFound model was pretrained via DINOv2 [20], with specific adaptations in the model initialization strategy and pretraining objective weight to enhance its effectiveness in the CXR domain.
+For downstream adaptation, we introduce a Global and Local Representations Integration (GLoRI) head that was trained on top of the frozen CheXFound model to perform attention probing.
+GLoRI employs the attention mechanism to compute fine- and coarse-grained disease-specific local features, which are then integrated with the global image features to improve multilabel classification performance.
+We assessed CheXFound’s performance on a wide range of CXR interpretation tasks, from thoracic disease detection, opportunistic risk estimation, and malpositioned tube detection to anatomical structure segmentation.
+CheXFound outperformed previous state-of-the-art models such as RAD-DINO [12], EVA-X [11], and CheXzero [13] for classifying 40 disease findings at different prevalence levels on the CXR-LT 24 dataset [21].
+Besides, CheXFound demonstrated superior label efficiency, achieving best-performing results on the Shenzhen, Montgomery, and JSRT datasets with limited training data.
+We also found that CheXFound achieved significant performance increases compared with its comparisons for the out-of-distribution tasks, including opportunistic CVD risk estimation and mortality prediction on the PLCO dataset [22], disease detection on the VinDr-CXR dataset [23], malpositioned tube detection on the RANZCR-CLiP dataset [24], and rib segmentation on the VinDr-RibCXR dataset [25].
 Overall, we demonstrate CheXFound’s strong generalization capabilities across a wide range of downstream tasks on in-distribution and out-of-distribution datasets.
-CheXFound’s strong representation quality can enable diverse downstream adaptations with improved label efficiency.
+CheXFound’s strong representation quality will enable diverse downstream adaptations with improved label efficiency in future applications.
 2 Related works
@@
 2.2 Foundation Models for Medical Applications
+Prior to the advent of foundation models, studies used both end-to-end supervised training and supervised fine-tuning methods to develop models for chest X-ray analysis.
+For example, CheXNet [30] and CheXpert [15] trained convolution neural networks from scratch, achieving better performance than radiologists for the detection of pneumonia and five selected pathologies, respectively.
+In contrast, supervised fine-tuning methods divided network training into two stages: pretraining to initialize network parameters and fine-tuning to adapt the network to specific tasks.
+Such methods include CheXtransfer [31], MedAug [32], and MoCo-CXR [33].
+CheXtransfer directly fine-tuned a convolutional neural network pretrained on ImageNet for CXR analysis, allowing the model to effectively take advantage of natural domain knowledge to achieve enhanced performance.
+Both MedAug and MoCo-CXR improved the paradigm of CheXtransfer by pretraining their own models on CXR images via contrastive learning before fine-tuning them on disease detection datasets.
 The surge in available data and computational resources have enabled the large-scale pretraining of foundation models.
 Studies have demonstrated that scaling foundation models in data and model sizes can achieve performance increases across a wide array of downstream tasks [7, 20, 28].
-In medical domain, research works have developed multiple categories of foundation models differing in technical approaches and data modalities.
-Our study is related to vision-centric foundation models.
-RAD-DINO [12] and EVA-X [11] are two foundation models in CXR domain.
-Compared to CheXFound with ViT-L pretrained on CXR-1M, these models are limited in model and data scales.
+In medical domain, studies have developed multiple categories of foundation models that use different pretraining approaches and data modalities.
+Our study belongs to the family of vision-centric foundation models.
+In the domain of CXR image analysis, several studies have explored various strategies for developing vision-centric foundation models, including RAD-DINO [12], RayDINO [34], EVA-X [11], Medical MAE [35], Foundation Ark [36], and Adam [37].
+However, as summarized in Table 1, most of these related studies remain limited in either model capacity or data scale.
+Models such as [12, 11, 35, 36, 37] typically employed the ViT-Base or Swin Transformer-Base architecture, with up to 88M parameters, and were pretrained on up to 838K CXRs from five public datasets.
+To further investigate the scalability, this work scales CheXFound to the ViT-Large architecture and pretrains it on 987K CXRs from 12 different sources.
+Among the above CXR foundation models, only RayDINO was based on the ViT-Large architecture, as adopted in this work.
+However, RayDINO did not sufficiently exploit the benefit of data diversity, relying on only four commonly used public datasets for pretraining.
+In contrast, our work expanded the pretraining dataset with eight additional sources, resulting in 124K more pretraining images.
+Table 1: Summary of the network architectures and pretraining datasets utilized by recent vision-centric foundation models.
+Models Architectures # Parameters # CXRs # Data sources Medical MAE [35] ViT-Base 86M 509K 3 public datasets Adam [37] ResNet-50 25M 224K 1 public datasets Foundation Ark [36] Swin Transformer-Base 88M 704K 4 public datasets EVA-X [11] ViT-Base 86M 520K 3 public datasets RAD-DINO [12] ViT-Base 86M 838K 5 public datasets + private dataset RayDINO [34] ViT-Large 307M 863K 4 public datasets CheXFound ViT-Large 307M 987K 12 public datasets
 Another category of foundation models incorporate vision and text data for multimodal pretraining.
@@
 Overall, research empirically finds that vision-language models achieve inferior performance than vision-centric foundation models in CXR classification [12].
-In this study, we focus on the vision-centric foundation model and investigate its capability for extensitve CXR classification tasks.
+In this study, we focus on the vision-centric foundation model and investigate its capability for extensive CXR classification tasks.
 To the best of our knowledge, our work employs the largest-scale self-supervised pretraining with over 1 million unique CXRs.
@@
 Figure 1: Overview of self-supervised pretraining of CheXFound, using publicly available CXRs from multiple institutions with a masked image modeling objective and a [CLS] token alignment objective.
-Table 1: Curation of CXR-1M with publicly available datasets from diverse institutions for self-supervised pretraining.
-CXR-1M is subsetted into CXR-207K and CXR-744K to evaluate the data scalability of self-supervised models.
+Table 2: Curation of CXR-987K with publicly available datasets from diverse institutions for self-supervised pretraining.
+CXR-987K is subsetted into CXR-207K and CXR-744K to evaluate the data scalability of self-supervised models.
 Datasets View Findings image #
@@
 PadChest[16] Frontal, Lateral 193 diseases 160,861
-CXR14[17] Frontal 14 diseases 112,120
-BRAX[18] Frontal, Lateal 14 diseases 40,967
+NIH-CXR[17] Frontal 14 diseases 112,120
+BRAX[18] Frontal, Laterall 14 diseases 40,967
 Total number of images in CXR-744K: 744,692
-VinDr-CXR[19] Frontal 28 diseases 18,000
 CANDID-PTX[19] Frontal Pneumothorax 19,237
@@
 BIMCV COVID+[48] Frontal, Lateral COVID-19 80,889
-Total number of images in CXR-1M: 1,005,733
-3.1 CXR-1M for Pretraining CheXFound
-As detailed in Table 1, we curated the CXR-1M dataset for self-supervised pretraining by retrieving in total 1,005,733 unique CXRs from 13 publicly available datasets [14, 15, 16, 17, 36, 18, 37, 38, 19, 35, 34, 20, 33] that were released for various downstream tasks, including disease diagnosis, abnormality detection, foreign objection detection, and segmentation.
-To learn comprehensive representations for multiview CXR analysis, both frontal-view in PA (posterior-anterior) or AP (anterior-posterior) and lateral-view CXRs were included into CXR-1M.
-To evaluate the data scalability of self-supervised pretraining, we further created CXR-207K and CXR-744K, denoting two subsets of CXR-1M, as shown in Table 1.
+Total number of images in CXR-987K: 987,733
+3.1 Pretraining CheXFound with Large-scale Dataset
+As detailed in Table 2, we curated the CXR-987K dataset for self-supervised pretraining by retrieving in total 987,733 unique CXRs from 12 publicly available datasets [14, 15, 16, 17, 46, 18, 47, 48, 23, 45, 44, 19, 43] that were released for various downstream tasks, including disease diagnosis, abnormality detection, foreign objection detection, and segmentation.
+To learn comprehensive representations for multiview CXR analysis, both frontal-view in PA (posterior-anterior) or AP (anterior-posterior) and lateral-view CXRs were included into CXR-987K.
+To evaluate the data scalability of self-supervised pretraining, we further created CXR-207K and CXR-744K, denoting two subsets of CXR-987K, as shown in Table 2.
 CXR-207K contains approximately 207K CXRs from MIMIC-CXR.
-CXR-744K contains around 744K CXRs from five datasets: MIMIC-CXR, CheXpert, PadChest, CXR14, and BRAX.
-We used DINOv2 [21], a state-of-the-art self-supervised learning method, to pretrain CheXFound on CXR-1M.
+CXR-744K contains around 744K CXRs from five datasets: MIMIC-CXR, CheXpert, PadChest, NIH-CXR, and BRAX.
+We used DINOv2 [20], a state-of-the-art self-supervised learning method, to pretrain CheXFound on CXR-987K.
 DINOv2 inherits designs from DINO [8] and iBOT [9] and incorporates two self-distillation objectives: the masked image modeling loss $\mathcal{L}_{\texttt{MIM}}$ and the [CLS] token alignment loss $\mathcal{L}_{\texttt{[CLS]}}$.
@@
 This approach aims to train the network to learn high-level class representations with off-the-shelf linear probe capabilities.
-3.2 Global and Local Representation Integration for Multilabel Classification
+3.2 Global and Local Representations Integration for Multilabel Classification
 For the downstream evaluation of CheXFound, the linear probe classifier is a pivotal tool to evaluate the quality of pretrained representations.
@@
 In contrast, patch tokens from our pretrained CheXFound contain rich CXR representations and high-level contextual information learned via masked image modeling, which can provide disease-specific local features to substantially reduce ambiguities arisen from using the [CLS] token for classifying multiple pathologies.
-To take advantage of both local and global features for disease classification, we introduce GLoRI (Fig.
-2), which utilizes a cross-attention layer with disease queries to summarize patch token features and a skip-connection to integrate the [CLS] token towards final prediction.
-To be specific, GLoRI receives output patch tokens $\mathbf{u}^{\texttt{Patch}}\in\mathbb{R}^{N\times D_{\text{model}}}$ from the frozen CheXFound backbone as input, where $D_{\text{model}}$ is the backbone embedding dimension.
+To take advantage of both local and global features for multilabel disease classification, we introduce GLoRI (Fig.
+2), which consists of two local-feature branches that use an adaptive temperature module to provide fine-grained local features and a pyramid patch merging module to provide coarse-grained local features respectively and a skip connection to integrate the global image features from the [CLS] token for multilabel disease classification.
+3.2.1 Encoding Fine-grained Local Features with Adaptive Temperatures
+GLoRI receives output patch tokens $\mathbf{u}^{\texttt{Patch}}\in\mathbb{R}^{N\times D_{\text{model}}}$ from the frozen CheXFound backbone as input, where $D_{\text{model}}$ is the backbone embedding dimension.
 Since there can be a dimension mismatch between the backbone and GLoRI, we use a linear embedding layer to project $\mathbf{u}^{\texttt{Patch}}$ to the GLoRI dimensional space:
@@
 where $\mathbf{u}^{\prime\texttt{Patch}}\in\mathbb{R}^{N\times D_{\text{GLoRI}}}$ is the projected by the linear embedding layer $\text{Linear}^{\text{embed}}(\cdot)$ to the patch token sequence with dimension $D_{\text{GLoRI}}$.
-In GLoRI, to extract disease-specific local features using the cross-attention layer, we initialize $M$ disease queries corresponding to $M$ disease findings, denoted as $\mathbf{q}\in\mathbb{R}^{M\times D_{\text{key}}}$.
-The keys $\mathbf{k}\in\mathbb{R}^{M\times D_{\text{key}}}$ and values $\mathbf{q}\in\mathbb{R}^{M\times D_{\text{value}}}$ for the cross-attention layer are from $\mathbf{u}^{\prime\texttt{Patch}}$, which provides rich CXR representations.
-A scaled dot-product attention module is used to compute attention-pooled features relevant to the query diseases:
-$\mathbf{q}^{\prime}=\text{Softmax}\left(\frac{\text{Linear}^{\text{query}}(\mathbf{q})\text{Linear}^{\text{key}}(\mathbf{k})^{T}}{\sqrt{D_{\text{key}}}}\right)\text{Linear}^{\text{value}}(\mathbf{v}),$ (2)
-where $\mathbf{q}^{\prime}\in\mathbb{R}^{M\times D_{\text{key}}}$ denotes the output disease queries, which we consider contain disease-specific local features.
-$D_{\text{key}}$ is the dimension of the key tokens.
-Last, we concatenate $\mathbf{q}^{\prime}$ with the [CLS] token to construct the GLoRI output token sequence.
-For multilabel classification, each GLoRI output token is projected by a linear classifier supervised by a binary cross-entropy loss.
-Overall, GLoRI applies the attentional principle, which shares similarity with previous works, such as Perceiver [39] and DETR [40], for CheXFound’s downstream evaluation.
-We empirically demonstrate that GLoRI extracts local features relevant to disease abnormalities in CXRs in Section 6.1.
+In GLoRI, to extract disease-specific local features using the multi-head attention layer, we initialize $M$ disease queries corresponding to $M$ disease findings, denoted as $\mathbf{q}\in\mathbb{R}^{M\times D_{\text{key}}}$.
+The keys $\mathbf{k}\in\mathbb{R}^{N\times D_{\text{key}}}$ and values $\mathbf{v}\in\mathbb{R}^{N\times D_{\text{value}}}$ for the multi-head attention layer are derived from $\mathbf{u}^{\prime\texttt{Patch}}$ to provide rich CXR representations.
+Additionally, to improve GLoRI’s capability to focus on small or diffuse abnormal regions, we introduce an adaptive temperature module to adjust the attention in the multi-head attention layer.
+The adaptive temperature module projects the average-pooled features of $\mathbf{u}^{\texttt{Patch}}$ through a multilayer perceptron with a tanh activation function and then applies a natural exponential function to generate the temperature vector $\boldsymbol{\uptau}\in\mathbb{R}^{M}$, where $M$ corresponds to the number of disease queries.
+We incorporate $\boldsymbol{\uptau}$ into the scaled dot-product attention mechanism to capture fine-grained local features:
+${\color[rgb]{0,0,0}\mathbf{q}^{\prime}=\text{Softmax}\left(\frac{\mathbf{q}\mathbf{k}^{T}}{\sqrt{D_{\text{key}}}\boldsymbol{\uptau}}\right)\mathbf{v},}$ (2)
+where $\mathbf{q}^{\prime}\in\mathbb{R}^{M\times D_{\text{key}}}$ denotes the output disease queries that capture fine-grained local features.
+$\boldsymbol{\uptau}$ adjusts the distribution of attention weights for each disease query to be either sharper or smoother with a value lower or larger than 1 respectively.
+3.2.2 Encoding Coarse-grained Local Features with Pyramid Patch Merging
+To effectively summarize contextual information, we incorporate a pyramid patch merging module to integrate multi-scale patch embeddings at low resolutions to provide coarse-grained local features.
+The pyramid patch merging module merges $8\times 8$, $4\times 4$, and $2\times 2$ adjacent patch tokens in $\mathbf{u}^{\texttt{Patch}}$ via average pooling to generate multi-scale patch embeddings.
+Then, it projects these patch embeddings to a dimension of $\frac{D_{\text{model}}}{3}$ via a linear layer and a ReLU activation, upsamples them to the original resolution, and concatenate them to construct patch embeddings with a dimension of $D_{\text{model}}$.
+We then apply a layer normalization to the patch embeddings before inputting them to the multi-head attention to encode coarse-grained local features.
+3.2.3 Integrating Global Image Features
+Last, we integrate the fine- and coarse-grained local features with the [CLS] token to construct the GLoRI output token sequence.
+For multilabel classification, each GLoRI output token represents a disease finding and is projected by a linear classifier, optimized using a binary cross-entropy loss, to perform disease detection.
+Overall, GLoRI employs the attention pooling mechanism, a strategy also employed by CLIP [49], V-JEPA [50], RayDINO [34], and GLoRIA [51] that incorporated attention pooling for image representation pooling [49, 51], video classification [50], and disease detection [34].
+GLoRI distinguishes itself from these related works by integrating two local-feature branches to capture both fine- and coarse-grained local features.
+We empirically demonstrate that GLoRI can extract local features relevant to disease abnormalities in CXRs in Section 6.1.
 Figure 2: Global and Local Representations Integration (GLoRI) for evaluating CheXFound on downstream tasks.
-GLoRI is appended on top of the frozen CheXFound backbone.
-GLoRI uses disease queries to compute attention-pooled CXR representations and integrates the global image features from [CLS] token for disease finding classification.
-For the opportunistic CXR interpretation, additional disease queries are created for GLoRI to retrieve representations related to the CVD risk estimation and mortality prediction.
+GLoRI is trained on top of the frozen CheXFound backbone.
+GLoRI incorporates adaptive temperatures and pyramid patch merging into the attention mechanism to encode fine-grained and coarse-grained local features respectively and integrates the global image features from the [CLS] token to predict final multilabel classification results.
 4 Experimental Design
@@
 4.1.1 Self-supervised pretraining details
-We conducted self-supervised pretraining of CheXFound on our curated CXR-1M dataset (Table 1) using the ViT-L architecture with a patch size of 16$\times$16 pixels.
+We conducted self-supervised pretraining of CheXFound on our curated CXR-987K dataset (Table 2) using the ViT-L architecture with a patch size of 16$\times$16 pixels.
 The loss weights for $\mathcal{L}_{\texttt{[CLS]}}$ and $\mathcal{L}_{\texttt{MIM}}$ were set to 1.0 and 3.0, respectively.
 The momentum to compute the exponential moving average of the student network was set to 0.994.
-We varied the global and local crop sizes to pretrain CheXFound at a rang of resolutions.
+We varied the global and local crop sizes to pretrain CheXFound at a range of resolutions.
 Specifically, we set the global and local crop size pairs to be (512, 144), (448, 128), (336, 128), and (224, 96).
@@
 Depending on the image resolutions, the pretraining processes take around 48 to 96 hours.
-The project source code is publicly available at https://github.com/RPIDIAL/CheXFound.
 4.1.2 Downstream evaluation details
@@
 We then combined the training and validation sets for a second round of training using the best learning rate.
-To maximize the number of images that the GLoRI module processes during downstream adaptation, we trained GLoRI for 10 epochs on CXR-LT 24, CheXpert, and PLCO and 100 epochs on Shenzhen, Mongomery, and JSRT.
+To maximize the number of images that the GLoRI head processes during downstream adaptation, we trained GLoRI for 10 epochs on CXR-LT 24, CheXpert, and PLCO and 100 epochs on Shenzhen, Mongomery, and JSRT.
 We set the batch size to 16 for Montgomery and JSRT and 256 for CXR-LT 24, CheXpert, Shenzhen, and PLCO.
-4.2 Experimental Design
-To rigorously evaluate CheXFound’s in-distribution and out-of-distribution performance, we employed an extensive classification benchmark, consisting of in-distribution datasets (CXR-LT 24 [22] and CheXpert [15]) and out-of-distribution datasets (Shenzhen [41], Montgomery[41], Japanese Society of Radiological Technology (JSRT) [42], and Prostate, Lung, Colorectal, and Ovarian (PLCO) Cancer Screening Trial [23]).
+4.2 Experimental datasets
+To rigorously evaluate CheXFound’s in-distribution and out-of-distribution performance, we employed an extensive classification benchmark, consisting of in-distribution datasets (CXR-LT 24 [21] and CheXpert [15]) and out-of-distribution datasets (VinDr-CXR [23], CXR-Pneumonia [52], Shenzhen [53], Montgomery[53], Japanese Society of Radiological Technology (JSRT) [54], RANZCR-CLiP [24], and Prostate, Lung, Colorectal, and Ovarian (PLCO) Cancer Screening Trial [22]).
 Following recent works in vision-centric foundation models [12, 28], we split the evaluation datasets into training, validation, and test splits.
 To avoid any potential data contamination, we included only the training set in CXR-LT 24 and the training and validation sets in CheXpert for self-supervised pretraining, while keeping the test set unseen.
-For the out-of-distribution datasets (Shenzhen, Montgomery, and JSRT), none of the images in the training, validation, and test sets were used for self-supervised pretraining.
+For the out-of-distribution datasets (VinDr-CXR, CXR-Pneumonia, Shenzhen, Montgomery, JSRT, RANZCR-CLiP, and PLCO), none of the images in the training, validation, and test sets were used for self-supervised pretraining.
 Although the MIMIC-CXR dataset is the common benchmark to evaluate the performance of CXR interpretation models, its labels contain only 14 findings.
@@
 It was divided into 191,027 frontal-view images in the training set, 202 images in the validation set, and 518 images in the test set.
-To assess the out-of-distribution generalization capabilities of the foundation models, we also performed evaluation on the Shenzhen and Montgomery datasets for tuberculosis detection and the JSRT database for lung nodule detection.
+To assess the out-of-distribution generalization capabilities of the foundation models, we evaluated model performance on VinDr-CXR that contains labels for 28 diseases.
+We used the official data splits of VinDr-CXR that contain 15000 CXRs for training and 3000 CXRs for test.
+We also performed evaluation on CXR-Pneumonia for pneumonia detection, Shenzhen and Montgomery for tuberculosis detection, and JSRT for lung nodule detection.
+The CXR-Pneumonia dataset contains 5,232 / 624 CXRs for training and test respectively.
 The Shenzhen, Montgomery, and JSRT datasets were divided into training, validation, and test splits with a ratio of 70:10:20.
@@
 JSRT contains 171 training images, 24 validation images, and 50 test images.
+In addition, we evaluated model performance for malpositioned tube detection on RANZCR-CLiP.
+We split the RANZCR-CLiP dataset into 18,742 / 2,678 / 5,356 CXRs for training, validation, and test respectively.
 To evaluate the extended predictive power of the foundation models, we obtained the lung screening CXRs from the PLCO trial and extracted the all-cause mortality and cardiovascular disease mortality labels from the up to 25-year follow-up data.
@@
 We estimated the 95% confidence intervals of the model performance in AUPRC and AUROC over 1,000 bootstrapped samples.
-To test statistical significance, we used a two-sided paired permutation test with 1,000 permutations to assess the observed performance differences of the two models for disease findings.
+To test statistical significance, we used a two-sided permutation test with 1,000 permutations to assess the observed performance differences of the two models for disease findings.
 5 Experimental Results
-Table 2: Comparison CheXFound and other foundation models when using linear probe and GLoRI for classification.
+Table 3: Classification results of CheXFound and other foundation models using linear probe and GLoRI on in-distribution datasets.
 Values inside the parentheses indicate the 95% confidence intervals.
 Values in bold indicate the best results.
-Classifier Foundation CXR-LT 24 CheXpert Shenzhen Montgomery JSRT
+Classifier Foundation CXR-LT 24 CheXpert
+Methods Models AUPRC AUROC AUPRC AUROC
+Linear probe PubMedCLIP [39] 0.089(.088–.089) 0.561(.554–.568) 0.277(.244–.311) 0.595(.564–.628)
+BiomedCLIP [38] 0.117(.116–.118) 0.643(.636–.649) 0.557(.504–.607) 0.841(.822–.860)
+CheXzero [13] 0.112(.111–.112) 0.552(.545–.558) 0.468(.424–.509) 0.778(.754–.798)
+EVA-X [11] 0.114(.113–.115) 0.596(.590–.602) 0.468(.422–.512) 0.788(.763–.812)
+RAD-DINO [12] 0.114(.113–.114) 0.557(.570–.583) 0.463(.422–.503) 0.746(.715–.778)
+CheXFound 0.209(.204–.214) 0.799(.794–.804) 0.620(.630–.727) 0.876(.860–.892)
+GLoRI head PubMedCLIP [39] 0.116(.115–.117) 0.649(.643–.655) 0.501(.450–.552) 0.804(.778–.828)
+BiomedCLIP [38] 0.122(.121–.123) 0.643(.636–.649) 0.552(.506–.593) 0.829(.809–.847)
+CheXzero [13] 0.131(.130–.132) 0.671(.665–.677) 0.599(.551–.647) 0.888(.868–.905)
+EVA-X [11] 0.149(.147–.150) 0.679(.672–.685) 0.614(.571–.659) 0.870(.853–.888)
+RAD-DINO [12] 0.173(.171–.176) 0.723(.717–.729) 0.639(.593–.687) 0.884(.869–.898)
+CheXFound 0.265(.259–.271) 0.840(.836–.844) 0.679(.630–.727) 0.908(.894–.921)
+Table 4: Classification results of CheXFound and other foundation models on out-of-distribution datasets.
+Values inside the parentheses indicate the 95% confidence intervals.
+Values in bold indicate the best results.
+Classifier Foundation VinDr-CXR CXR-Pneumonia Shenzhen Montgomery JSRT
 Methods Models AUPRC AUROC AUPRC AUROC AUPRC AUROC AUPRC AUROC AUPRC AUROC
-Linear probe PubMedCLIP [29] 0.089(.088–.089) 0.561(.554–.568) 0.277(.244–.311) 0.595(.564–.628) 0.857(.786–.913) 0.814(.738–.887) 0.565(.306–.817) 0.534(.310–.750) 0.785(.630–.912) 0.685(.511–.833)
-BiomedCLIP [28] 0.117(.116–.118) 0.643(.636–.649) 0.557(.504–.607) 0.841(.822–.860) 0.903(.843–.949) 0.885(.827–.934) 0.925(.791–1.0) 0.929(.822–1.0) 0.589(.427–.755) 0.432(.265–.601)
-CheXzero [13] 0.112(.111–.112) 0.552(.545–.558) 0.468(.424–.509) 0.778(.754–.798) 0.929(.884–.962) 0.906(.851–.950) 0.964(.869–1.0) 0.970(.898–1.0) 0.759(.597–.885) 0.620(.453–.777)
-EVA-X [11] 0.114(.113–.115) 0.596(.590–.602) 0.468(.422–.512) 0.788(.763–.812) 0.840(.738–.921) 0.824(.746–.898) 0.577(.320–.809) 0.509(.278–.749) 0.641(.472–.797) 0.490(.312–.660)
-RAD-DINO [12] 0.114(.113–.114) 0.557(.570–.583) 0.463(.422–.503) 0.746(.715–.778) 0.883(.818–.933) 0.861(.798–.916) 0.637(.369–.842) 0.561(.316–.788) 0.747(.585–.888) 0.623(.456–.778)
-CheXFound 0.209(.204–.214) 0.799(.794–.804) 0.620(.630–.727) 0.876(.860–.892) 0.974(.949–.992) 0.967(.935–.990) 0.988(.939–1.0) 0.990(.952–1.0) 0.918(.826–.975) 0.856(.741–.948)
-GLoRI head PubMedCLIP [29] 0.116(.115–.117) 0.649(.643–.655) 0.501(.450–.552) 0.804(.778–.828) 0.897(.832–.946) 0.867(.807–.927) 0.628(.363–.866) 0.694(.484–.872) 0.652(.483–.810) 0.493(.330–.654)
-BiomedCLIP [28] 0.122(.121–.123) 0.643(.636–.649) 0.552(.506–.593) 0.829(.809–.847) 0.921(.870–.957) 0.897(.840–.944) 0.900(.738–1.0) 0.898(.744–1.0) 0.634(.470–.789) 0.505(.291–.704)
-CheXzero [13] 0.131(.130–.132) 0.671(.665–.677) 0.599(.551–.647) 0.888(.868–.905) 0.912(.852–.954) 0.894(.838–.942) 0.726(.451–.901) 0.653(.396–.882) 0.640(.467–.807) 0.485(.323–.662)
-EVA-X [11] 0.149(.147–.150) 0.679(.672–.685) 0.614(.571–.659) 0.870(.853–.888) 0.928(.881–.964) 0.896(.835–.945) 0.977(.909–1.0) 0.979(.918–1.0) 0.866(.760–.936) 0.748(.615–.864)
-RAD-DINO [12] 0.173(.171–.176) 0.723(.717–.729) 0.639(.593–.687) 0.884(.869–.898) 0.909(.854–.952) 0.885(.823–.936) 0.909(.747–1.0) 0.911(.782–1.0) 0.683(.503–.861) 0.615(.440–.783)
-CheXFound 0.252(.247–.258) 0.830(.826–.834) 0.679(.630–.727) 0.908(.894–.921) 0.983(.960–.996) 0.978(.951–.995) 1.000(1.0–1.0) 1.000(1.0–1.0) 0.986(.956–1.0) 0.975(.931–1.0)
+Linear probe PubMedCLIP [39] 0.088(.082–.095) 0.570(.514–.609) 0.954(.936–.969) 0.924(.904–.944) 0.857(.786–.913) 0.814(.738–.887) 0.565(.306–.817) 0.534(.310–.750) 0.785(.630–.912) 0.685(.511–.833)
+BiomedCLIP [38] 0.100(.094–.106) 0.587(.544–.618) 0.984(.976–.990) 0.973(.962–.986) 0.903(.843–.949) 0.885(.827–.934) 0.925(.791–1.0) 0.929(.822–1.0) 0.589(.427–.755) 0.432(.265–.601)
+CheXzero [13] 0.102(.096–.109) 0.536(.486–.567) 0.973(.962–.981) 0.953(.938–.967) 0.929(.884–.962) 0.906(.851–.950) 0.964(.869–1.0) 0.970(.898–1.0) 0.759(.597–.885) 0.620(.453–.777)
+EVA-X [11] 0.105(.099–.111) 0.542(.503–.572) 0.975(.959–.988) 0.968(.953–.980) 0.840(.738–.921) 0.824(.746–.898) 0.577(.320–.809) 0.509(.278–.749) 0.641(.472–.797) 0.490(.312–.660)
+RAD-DINO [12] 0.102(.096–.107) 0.477(.441–.507) 0.995(.991–.998) 0.992(.986–.996) 0.883(.818–.933) 0.861(.798–.916) 0.637(.369–.842) 0.561(.316–.788) 0.747(.585–.888) 0.623(.456–.778)
+CheXFound 0.298(.280–.317) 0.838(.776–.866) 0.996(.993–.998) 0.993(.988–.997) 0.974(.949–.992) 0.967(.935–.990) 0.988(.939–1.0) 0.990(.952–1.0) 0.918(.826–.975) 0.856(.741–.948)
+GLoRI head PubMedCLIP [39] 0.104(.098–.113) 0.589(.532–.640) 0.973(.961–.983) 0.960(.944–.973) 0.897(.832–.946) 0.867(.807–.927) 0.628(.363–.866) 0.694(.484–.872) 0.652(.483–.810) 0.493(.330–.654)
+BiomedCLIP [38] 0.094(.089–.102) 0.537(.497–.566) 0.987(.981–.992) 0.978(.969–.986) 0.921(.870–.957) 0.897(.840–.944) 0.900(.738–1.0) 0.898(.744–1.0) 0.634(.470–.789) 0.505(.291–.704)
+CheXzero [13] 0.074(.072–.078) 0.489(.446–.515) 0.982(.972–.989) 0.972(.959–.982) 0.912(.852–.954) 0.894(.838–.942) 0.726(.451–.901) 0.653(.396–.882) 0.640(.467–.807) 0.485(.323–.662)
+EVA-X [11] 0.163(.155–.172) 0.662(.610–.694) 0.990(.979–.997) 0.988(.978–.995) 0.928(.881–.964) 0.896(.835–.945) 0.977(.909–1.0) 0.979(.918–1.0) 0.866(.760–.936) 0.748(.615–.864)
+RAD-DINO [12] 0.155(.147–.163) 0.680(.630–.718) 0.994(.991–.997) 0.991(.985–.995) 0.909(.854–.952) 0.885(.823–.936) 0.909(.747–1.0) 0.911(.782–1.0) 0.683(.503–.861) 0.615(.440–.783)
+CheXFound 0.341(.319–.367) 0.872(.813–.898) 0.996(.992–.998) 0.993(.987–.997) 0.983(.960–.996) 0.978(.951–.995) 1.000(1.0–1.0) 1.000(1.0–1.0) 0.986(.956–1.0) 0.975(.931–1.0)
 Table 5: Comparison of CheXFound with the end-to-end trained model and vision-language foundation models.
@@
 CheXzero [13] 0.134(.133–.136) 0.668(.662–.674) 0.646(.600–.692) 0.888(.868–.905) 0.875(.804–.933) 0.849(.776–.911) 0.967(.870–1.0) 0.969(.889–1.0) 0.708(.532–.863) 0.530(.378–.698)
-GLoRI head CheXFound 0.252(.247–.258) 0.830(.826–.834) 0.679(.630–.727) 0.908(.894–.921) 0.983(.960–.996) 0.978(.951–.995) 1.000(1.0–1.0) 1.000(1.0–1.0) 0.986(.956–1.0) 0.975(.931–1.0)
+GLoRI head CheXFound 0.265(.259–.271) 0.840(.836–.844) 0.679(.630–.727) 0.908(.894–.921) 0.983(.960–.996) 0.978(.951–.995) 1.000(1.0–1.0) 1.000(1.0–1.0) 0.986(.956–1.0) 0.975(.931–1.0)
 5.1 Overall performance comparison
 A pivotal characteristic of foundation models lies in their capability to achieve improved performance on a wide range of downstream datasets.
-To evaluate the capability of foundation models, we compared CheXFound, which uses ViT-L pretrained on CXR-1M, with publicly available pretrained encoders, including RAD-DINO [12], EVA-X [11], CheXzero [13], BiomedCLIP [28], PubmedCLIP [29], and ConvNeXt [43].
-RAD-DINO was pretrained on a combined dataset comprising MIMIC-CXR, CheXpert, PadChest, CXR14, and BRAX, using the DINOv2 framework.
-EVA-X was pretrained on MIMIC-CXR, CheXpert, and CXR14 using EVA [44] technique with contrastive vision features for masked image modeling.
+To evaluate the capability of foundation models, we compared CheXFound, which uses ViT-L pretrained on CXR-987K, with publicly available pretrained encoders, including RAD-DINO [12], EVA-X [11], CheXzero [13], BiomedCLIP [38], PubmedCLIP [39], and ConvNeXt [55].
+RAD-DINO was pretrained on a combined dataset comprising MIMIC-CXR, CheXpert, PadChest, NIH-CXR, and BRAX, using the DINOv2 framework.
+EVA-X was pretrained on MIMIC-CXR, CheXpert, and NIH-CXR using EVA [56] technique with contrastive vision features for masked image modeling.
 CheXzero, a vision-language foundation model, was pretrained on MIMIC-CXR using contrastive language-image pretraining (CLIP) [49].
@@
 PubmedCLIP was initialized with the CLIP model and finetuned on PubMed articles.
-Finally, ConvNeXt was pretraiend on the ImageNet-22K dataset.
+Finally, ConvNeXt was pretrained on the ImageNet-22K dataset.
 Figure 3: Detailed Performance for 40 disease findings in AUROC on the CXR-LT 24 dataset.
@@
 Error bars indicate the 95% confidence intervals of AUPRC over 1,000 bootstrapped samples.
-To evaluate the effectiveness and generalizability of the representations extracted by the foundation models, we evaluated the linear probe performance across the five datasets (CXR-LT 24, CheXpert, Shenzhen, Montgomery, and JSRT).
-While linear probe provides a straightforward approach to evaluate the quality of representation, it primarily relies on global image features from the [CLS] token, often resulting in suboptimal performance.
-Hence, we also evaluated the pretrained encoders using GLoRI, which incorporates attention-pooled local features in addition to the global image features from the [CLS] token.
-For vision-language pretained encoders, we further validated the quality of their vision representations by examing their correlation with text features for disease finding classification.
-Across all the five classification tasks (CXR-LT 24, CheXpert, Shenzhen, Montgomery, and JSRT), CheXFound consistently outperformed other foundation models in the linear probe setting as shown in Table 2.
-On the multilabel, long-tailed classification task (CXT-LT 24), CheXFound achieved an AUPRC of 0.209, outperforming the next best-performing model (either RAD-DINO or EVA-X) by 9.5% ($p<0.001$, two-sided paired permutation test).
-On the five-class multilabel classification task (CheXpert), CheXFound outperformed the next best-performing model (EVA-X) by 8.8% ($p<0.001$) in AUROC.
-On single-class classification tasks (Shenzhen, Montgomery, and JSRT) with limited amounts of training data, CheXFound similarly outperformed the next best-performing models in AUROC by 11.7% ($p<0.001$), 42.9% ($p<0.001$), and 23.3% ($p<0.001$), respectively.
-We further evaluated the performance of the foundation models using GLoRI across five classification tasks (Table 2).
+To evaluate the effectiveness and generalizability of the representations extracted by the foundation models, we evaluated the linear probe performance across seven datasets (CXR-LT 24, CheXpert, VinDr-CXR, CXR-Pneumonia, Shenzhen, Montgomery, and JSRT).
+While linear probe stands for a straightforward approach to evaluate the quality of representations, it primarily relies on global image features from the [CLS] token, often resulting in suboptimal performance.
+We hence evaluated the pretrained encoders using our proposed GLoRI, which incorporates attention-pooled local features in addition to the global image features to enhance multilabel classification.
+Regarding vision-language pretrained encoders, we further validated their vision representation quality by assessing their correlation with text features to perform zero-shot classification.
+Across all the in-distribution and out-of-distribution classification tasks (CXR-LT 24, CheXpert, VinDr-CXR, CXR-Pneumonia, Shenzhen, Montgomery, and JSRT), CheXFound consistently outperformed other foundation models in the setting using a simple linear probe classifier, as shown in Tables 3 and 4.
+On the multilabel, long-tailed classification task (CXT-LT 24), CheXFound achieved an AUPRC of 0.209, outperforming the next best-performing model (either RAD-DINO or EVA-X) by 9.5% ($p<0.001$, two-sided permutation test).
+On the five-class multilabel classification task (CheXpert), CheXFound outperformed the next best-performing model (BiomedCLIP) by 3.5% ($p<0.001$) in AUROC.
+On the out-of-distribution multilabel classification task (VinDr-CXR), CheXFound outperformed the next best-performing model (BiomedCLIP) by 25.1% ($p<0.001$) in AUROC.
+On out-of-distribution single-class classification tasks (CXR-Pneumonia, Shenzhen, Montgomery, and JSRT) with limited amounts of training data, CheXFound outperformed the next best-performing models in AUROC by 0.1% ($p>0.05$), 6.1% ($p<0.001$), 2.0% ($p>0.05$), and 17.1% ($p<0.001$), respectively.
+We further evaluated the performance of the foundation models using GLoRI across the seven classification tasks (Tables 3 and 4).
 The foundation models (CheXFound, RAD-DINO, EVA-X, CheXzero, BiomedCLIP, and PubMedCLIP) with GLoRI generally outperformed their linear probe baselines.
-Specifically, CheXFound with GLoRI outperformed its linear probe baseline in AUROC by 3.1%, 3.2%, 1.1%, 1.0%, and 11.9% on CXR-LT 24, CheXpert, Shenzhen, Montgomery, and JSRT, respectively.
-CheXFound with GLoRI also outperformed other foundation models, including CLIP-based models (CheXzero, BioMedCLIP, and PubMedCLIP) under image-text alignment, as well as end-to-end trained ConvNeXt (Table 3).
+CheXFound with GLoRI outperformed its linear probe baseline in AUROC by 4.1%, 3.2%, 3.4%, 1.1%, 1.0%, and 11.9% on CXR-LT 24, CheXpert, VinDr-CXR, Shenzhen, Montgomery, and JSRT, respectively.
+CheXFound achieved an AUROC of 0.993 on CXR-Pneumonia, on par with its linear probe baseline.
+In addition, as shown in Table 5, CheXFound with GLoRI outperformed CLIP-based foundation models (CheXzero, BioMedCLIP, and PubMedCLIP) when image-text alignment was used for zero-shot classification.
+CheXFound also outperformed the end-to-end trained model ConvNeXt.
 To show the detailed performance of CheXFound over 40 disease findings, we illustrated CheXFound performance in AUROC against its comparisons in Fig.
@@
 We also compared performance in AUPRC over disease findings with high, medium, and low prevalence in Fig.
-3b.
+4.
 CheXFound outperformed its comparisons under all levels of prevalence, even for underrepresented pathologies in the low prevalence category.
+5.2 Malpositioned tube detection
+The RANZCX-CLiP dataset contains CXRs for verifying the position of lines and tubes.
+This dataset categorized these CXRs into normal, borderline, and abnormal classes based on the position of tubes (endotracheal tube, nasogastric tube, and central venous catheter) to indicate whether they were appropriately positioned.
+The percentages of normal, borderline, and abnormal samples are 59.4%, 28.8%, and 11.8% respectively, representing an unbalanced class distribution.
+As shown in Table 6, CheXFound achieved AUPRCs of 0.857, 0.529, and 0.445 for classifying normal, borderline, and abnormal tubes respectively.
+CheXFound outperformed the next best-performing models by 16.4% ($p<0.001$), 16.1% ($p<0.001$), and 29.2% ($p<0.001$) for classifying normal, borderline, and abnormal classes respectively.
+These performance gains could be attributed to CheXFound’s capability to encode high-resolution CXRs and produce effective representations of nuanced structures.
+Table 6: Classification results of foundation models for malpositioned tube detection in CXRs.
+We computed the mean AUPRC values and their 95% confidence intervals over 1,000 bootstrapped samples.
+Values in bold indicate the best-performing results.
+Methods Normal Borderline Abnormal
+PubMedCLIP [39] 0.661(.642–.679) 0.348(.328–.372) 0.136(.123–.149)
+BiomedCLIP [38] 0.645(.626–.664) 0.330(.311–.350) 0.133(.121–.148)
+CheXzero [13] 0.693(.674–.711) 0.364(.341–.386) 0.153(.139–.173)
+EVA-X [20] 0.657(.637–.676) 0.339(.317–.361) 0.135(.122–.150)
+RAD-DINO [20] 0.690(.670–.709) 0.368(.347–.391) 0.138(.125–.154)
+CheXFound 0.857(.844–.869) 0.529(.503–.556) 0.445(.403–.488)
 5.3 Opportunistic predictive power
-Beyond thoracic disease detection tasks, we evaluated ChesXFound’s generalizability in opportunistic CXR interpretation.
-For this purpose, we requested access to the CXR arm of the PLCO trial [46], which includes digitally scanned CXR films and up to 25-year morality follow-up data.
+Beyond thoracic disease detection tasks, we evaluated CheXFound’s generalizability in opportunistic CXR interpretation.
+For this purpose, we requested access to the CXR arm of the PLCO trial [57], which includes digitally scanned CXR films and up to 25-year mortality follow-up data.
 Using this dataset, we investigated CheXFound’s predictive capability for cardiovascular disease (CVD) risk and all-cause mortality estimation.
-We used CheXFound with ViT-L pretrained on CXR-1M in this experiment and compared CheXFound against two vision foundation models (RAD-DINO, EVA-X) and the end-to-end trained model, ConvNeXt.
+We used CheXFound with ViT-L pretrained on CXR-987K in this experiment and compared CheXFound against two vision foundation models (RAD-DINO, EVA-X) and the end-to-end trained model, ConvNeXt.
 CheXFound consistently outperformed its counterparts in both CVD risk and all-cause mortality estimation tasks.
@@
 Shaded areas indicate 95% confidence intervals.
+5.4 CXR foundation models for semantic segmentation
+In addition to the classification tasks, we applied the pretrained CheXFound for semantic segmentation on CXRs.
+We included the VinDr-RibCXR [25] dataset which contains segmentation masks for 10 left ribs and 10 right ribs and the Montgomery dataset [53] which contains segmentation masks for left and right lungs.
+We trained a UPerNet decoder [58] on top of the frozen foundation models to perform segmentation.
+We compared CheXFound with other vision foundation models (EVA-X [11] and RAD-DINO [12]) and an end-to-end trained model (ConvNeXt [55]).
+As shown in Table 8, CheXFound achieved a Dice similarity coefficient of 0.793 on VinDr-RibCXR, outperforming the next best-performing model (ConvNeXt) by 7.1% ($p<0.001$).
+On Montgomery, CheXFound achieved a Dice similarity coefficient of 0.980, slightly surpassing the next best-performing model (RAD-DINO) by 0.2%.
+Such marginal performance gain over RAD-DINO on Montgomery likely reflects the relative simplicity of lung segmentation on this dataset, where all comparison methods achieved satisfactory performance.
+In contrast, the significant performance improvements in individual rib segmentation on VinDr-RibCXR demonstrates CheXFound’s superior capability to adapt to complicated anatomical structure segmentation tasks.
+Table 8: Semantic segmentation results of the vision foundation models and end-to-end trained model.
+We trained a UPerNet [58] decoder on top of the frozen foundation model.
+We computed the mean and standard deviation of Dice similarity coefficients to evaluate segmentation performance.
+Values in bold indicate the best-performing results.
+Models Frozen Backbone VinDr-RibCXR Montgomery
+ConvNeXt [55] 0.710 $\pm$ 0.116 0.954 $\pm$ 0.047
+EVA-X [11] 0.471 $\pm$ 0.108 0.956 $\pm$ 0.036
+RAD-DINO [12] 0.684 $\pm$ 0.094 0.978 $\pm$ 0.006
+CheXFound 0.793 $\pm$ 0.062 0.980 $\pm$ 0.005
+Table 9: Benchmarking foundation models pretrained using different self-supervised methods on CXR-987K.
+We evaluated the downstream classification performance on CXR-LT 24.
+Values inside the parentheses are 95% confidence intervals.
+Values in bold indicate the best-performing results.
+Methods Architectures AUPRC AUROC
+Medical MAE [35] ViT-B/16 0.193(.189–.197) 0.772(.767–.779)
+iBOT [9] ViT-L/16 0.212(.207–.217) 0.801(.796–.806)
+EVA-X [11] ViT-B/16 0.168(.165–.171) 0.703(.698–.708)
+RAD-DINO [12] ViT-B/14 0.185(.181–.189) 0.751(0.743–.759)
+CheXFound ViT-L/16 0.265(.259–.271) 0.840(.836–.844)
 Table 10: Model Performance across pretraining data sizes and model scales.
@@
 ViT-Large 0.217(.214–.221) 0.813(.801–.822) 0.643(.596–.692) 0.887(.876–.896) 0.964(.935–.986)) 0.956(.923–.972) 0.953(.843–1.0) 0.957(.845–1.0) 0.905(.794–.953) 0.826(.674–.925)
-CXR-1M ViT-Base 0.219(.215–.223) 0.815(.811–.819) 0.632(.583–.679) 0.877(.853–.896) 0.957(.924–.979) 0.947(.925–.963) 0.923(.788–1.0) 0.927(.818–1.0) 0.908(.806–.955) 0.845(.696–.943)
-ViT-Large 0.252(.247–.258) 0.830(.826–.834) 0.679(.630–.727) 0.908(.894–.921) 0.983(.960–.996) 0.978(.951–.995) 1.000(1.0–1.0) 1.000(1.0–1.0) 0.986(.956–1.0) 0.975(.931–1.0)
-5.3 Scalability of self-supervised vision encoders
-The scaling capabilities of the self-supervised vision encoders depends on both the model size and the pretraining data size and diversity [21, 26].
-To analyze the scaling trends, we pretrained CheXFound across a range of data scales, with CXR-1M and its two subsets CXR-744K and CXR-207K.
-We also evaluated the impact of model scale by using ViT-Base (ViT-B) and ViT-Large (ViT-L) as the backbones.
-Our results in Table 5 demonstrate that CheXFound benefits form both data and model scaling.
-Increasing the pretraining data from CXR-207K to CXR-1M with a ViT-L backbone leads to significant AUROC improvements of 3.5% ($p<0.001$) on CXR-LT 24, 3.4% ($p<0.001$) on CheXpert, 3.8% ($p<0.001$) on Shenzhen, 6.8% ($p<0.001$) on Montgomery, and 30.7% ($p<0.001$) on JSRT.
-We observed similar trends when using ViT-B, with performance also improving as we scale up the pretraining data from CXR-207K to CXR-1M.
-In addition, CheXFound with ViT-L consistently outperformed the ViT-B architecture across different data sizes.
-These results align with previous studies on scaling ViT models [21, 12, 26].
+CXR-987K ViT-Base 0.219(.215–.223) 0.815(.811–.819) 0.632(.583–.679) 0.877(.853–.896) 0.957(.924–.979) 0.947(.925–.963) 0.923(.788–1.0) 0.927(.818–1.0) 0.908(.806–.955) 0.845(.696–.943)
+ViT-Large 0.265(.259–.271) 0.840(.836–.844) 0.679(.630–.727) 0.908(.894–.921) 0.983(.960–.996) 0.978(.951–.995) 1.000(1.0–1.0) 1.000(1.0–1.0) 0.986(.956–1.0) 0.975(.931–1.0)
 Figure 6: Evaluation of CheXFound model on the CXR-LT 24, CheXpert, Shenzhen, Montgomery, and JSRT datasets across a range of pretraining image resolutions.
 Error bars indicate 95% confidence intervals.
-Figure 7: GLoRI attention maps for disease findings of atelectasis, cardiomegaly, consolidation, edema, and pleural effusion, respectively.
-Each subfigure contains 25 CXRs with their GLoRI attention maps overlaid (bottom) and an anchor CXR with a global attention map overlaid (top).
-The global attention maps are the averaged attention maps of 25 CXRs after registered to the anchor image.
+5.5 Benchmarking self-supervised pretraining methods
+We benckmarked various self-supervised pretraining methods on our curated CXR-987K dataset to evaluate their effectiveness in representation learning.
+By consistently using the CXR-987K dataset, we eliminated the influence of data heterogeneity in the pretraining stage.
+The self-supervised methods for comparison include Medical MAE [35], iBOT [9], EVA-X [11], and RAD-DINO [12].
+Classification performance was evaluated on the CXR-LT 24 dataset.
+As shown in Table 9 below, CheXFound outperformed the next best-performing method (iBOT) by 3.9% ($p<0.001$) in AUROC.
+CheXFound achieved higher performance than iBOT owing to our advanced pretraining approach that utilized a low-resolution warm-up strategy and an optimized MIM loss weight on top of the original DINOv2.
+Medical MAE underperformed CheXFound because it lacks a discriminative objective which is critical for classification tasks.
+RAD-DINO, EVA-X, and Medical MAE were constrained by the shallower ViT-Base architecture, which encoded less effective representations than the ViT-Large architecture employed in our work.
+5.6 Scalability of self-supervised vision encoders
+The capabilities of the self-supervised vision encoders are affected by both the pretraining data scale and model scale [20, 28].
+To analyze the scaling trends, we pretrained CheXFound across a range of data scales, including CXR-987K and its two subsets CXR-744K and CXR-207K.
+We also evaluated the impact of the model scale by using ViT-Base (ViT-B) and ViT-Large (ViT-L) as the backbones.
+We maintained consistent hyperparameters across all pretraining data sizes and model capacities.
+We trained CheXFound for 100 epochs (2,500 iterations per epoch) using the AdamW optimizer with a batch size of 14 and an initial learning rate of 2e-4.
+The resolution of input chest X-rays was set to 512$\times$512.
+We also applied a low-resolution warm-up strategy and a masked image modeling loss weight of 3, as determined by the ablation studies in Section 5.8.
+Our results in Table 10 demonstrate that CheXFound benefits form both data and model scaling.
+Increasing the pretraining data from CXR-207K to CXR-987K with the ViT-L backbone leads to significant AUROC improvements of 3.5% ($p<0.001$) on CXR-LT 24, 3.4% ($p<0.001$) on CheXpert, 3.8% ($p<0.001$) on Shenzhen, 6.8% ($p<0.001$) on Montgomery, and 30.7% ($p<0.001$) on JSRT.
+We observed similar trends when using ViT-B, which showed the performance consistently improved as we increased the pretraining data scale from CXR-207K to CXR-987K.
+In addition, CheXFound with ViT-L consistently outperformed the ViT-B architecture across different pretraining data scales.
+These results align with previous studies on scaling ViT models [20, 12, 28].
 5.7 Impact of CXR resolution
@@
 Increasing the pretraining resolutions from 2242 pixels to 5122 pixels significantly improves the AUROC by 3.3% ($p<0.001$), 3.5% ($p<0.001$), 3.7% ($p<0.001$), 10.7% ($p<0.001$), and 35.8% ($p<0.001$) on the CXR-LT 24, CheXpert, Shenzhen, Montgomery, and JSRT datasets, respectively.
+5.8 Ablation studies on pretraining
+We evaluated the impact of various parameter initialization strategies on downstream performance.
+As shown in Fig.
+7, random initialization and general domain initialization that utilized a checkpoint pretrained on natural images achieved AUROCs of 0.740 and 0.796, respectively, for classifying 40 diseases on CXR-LT 24.
+Incorporating a low-resolution warm-up strategy, which initializes pretraining at 512$\times$512 resolution with the checkpoint pretrained at 224$\times$224 resolution, increased the AUROC to 0.840, outperforming both the random initialization and general domain initialization strategies by large margins.
+We hypothesize that the degraded performance of direct high-resolution pretraining was due to the limited capability of the ViT architecture to capture extensive contextual information, which was mitigated by warming up the model with the low-resolution pretrained model that had already learned high-level contextual representations.
+We further evaluated the impact of pretraining loss weights.
+The original DINOv2 settings employ balanced weights of 1:1 for the [CLS] token alignment and MIM losses for representation learning with natural images, where local patches often contain sufficient information to infer the global view.
+In contrast, medical images are characterized by subtle differences, with diagnostic cues often residing in nuanced local features.
+As such, we conducted multiple runs of pretraining with increased MIM loss weights to search for an optimal weight factor that adequately encourages local feature learning.
+As illustrated in Fig.
+7, increasing MIM loss weights from 1 to 3 improved the AUROC from 0.822 to 0.840.
+Overall, compared with the original DINOv2 settings that achieved an AUROC of 0.793 (Fig.
+7) by training the model from scratch at 224$\times$224 resolution with a MIM loss weight of 1, our modifications to the initialization strategy for high-resolution pretraining and MIM loss weight delivered a performance gain of 4.7%.
+Figure 7: Ablation studies on the impact of MIM loss weights and parameter initialization strategies on downstream classification.
+We evaluated the classification performance on the CXR-LT 24 dataset.
+Error bars indicate 95% confidence intervals.
+Table 11: Ablation studies on the GLoRI architecture.
+We evaluated the classification performance on CXR-LT 24.
+We computed the mean AUPRC and AUROC values and their 95% confidence intervals over 1,000 bootstrapped samples.
+Values in bold indicate the best-performing results.
+Architecture AUPRC $p$-value AUROC $p$-value
+Linear probe 0.209(.204–.214) - 0.799(.794–.804) -
+Attention pooler 0.246(.241–.252) - 0.826(.822–.830) -
++ Global representation 0.252(.247–.258) 0.023 0.830(.826–.834) 0.021
++ Adaptive Temperatures 0.262(.256–.268) $<0.001$ 0.835(.831–.839) $<0.001$
++ Pyramid Patch Merging 0.265(.259–.271) $<0.001$ 0.840(.836–.844) $<0.001$
+Figure 8: Detailed results comparing GLoRI and linear probing for classifying different diseases.
+Diseases are sorted in descending order regarding their frequency on the dataset.
+We computed the mean and 95% confidence intervals of average precision using 1,000 bootstrapped samples.
+Error bars indicate 95% confidence intervals.
+Asterisks indicate statistical significance ($p<0.05$).”
+Figure 9: GLoRI attention maps for disease findings of atelectasis, cardiomegaly, consolidation, edema, and pleural effusion, respectively.
+Each subfigure contains 25 CXRs with their GLoRI attention maps overlaid (bottom) and an anchor CXR with a global attention map overlaid (top).
+The global attention maps are the averaged attention maps of 25 CXRs after registered to the anchor image.
+5.9 Ablation studies on GLoRI head
+We proposed GLoRI, a prediction head that incorporates both global and local image features for multilabel disease classification.
+Compared with prior attention pooling approaches [34, 50] that rely solely on an attention pooler with multi-head attention, GLoRI enhances classification performance by integrating an adaptive temperature module and a pyramid patch merging module into the multi-head attention mechanism and by incorporating the global image features.
+We evaluated GLoRI on the CXR-LT 24 dataset.
+As shown in Table 11, the attention pooler with a single multi-head attention layer achieved an AUROC of 0.826.
+This performance was improved by 0.4%, 0.5%, and 0.5% by incrementally incorporating the global representation, adaptive temperatures, and pyramid patch merging, respectively.
+Eventually, GLoRI achieved an AUROC of 0.84, which significantly improved the attention pooler by 1.4% ($p<0.001$, two-sided permutation test) and outperformed linear probe by 4.1% ($p<0.001$).
+We compared GLoRI with linear probe on detecting 40 diseases in Fig.
+8.
+Overall, these improvements are attributed to GLoRI’s enriched representations, which integrate both the fine-grained and coarse-grained local features alongside the global image features.
 6 Discussion and Conclusion
 6.1 Interpretation of disease-specific local features
-The interpretabliliy of an artificial intelligence model is crucial to its medical applications.
+The interpretability of an artificial intelligence model is crucial to its medical applications.
 In this study, we trained GLoRI on top of the frozen foundation model.
@@
 to provide a global perspective in Fig.
-7.
+9.
 The attention maps contain precise localization of abnormalities in CXRs and the global attention maps cover the regions where the pathologies constantly occur.
@@
 This pattern is well captured by our attention maps as we observed in Fig.
-7d that the critical regions in the individual attention maps scatter over the lungs and the global attention map covers extensive regions of both lungs.
+9 (d) that the critical regions in the individual attention maps scatter over the lungs and the global attention map covers extensive regions of both lungs.
 However, these attention maps have the limitations of covering partial regions of the abnormalities.
@@
 We also demonstrated CheXFound’s generalization capabilities on infrequent and underrepresented pathologies.
-CheXFound achieved significant increases over its comparions in classifying low-prevalence pathologies with lower than 1% occurrence frequencies, demonstrating its superior label efficiency.
+CheXFound achieved significant increases over its comparisons in classifying low-prevalence pathologies with lower than 1% occurrence frequencies, demonstrating its superior label efficiency.
 Although CheXFound with ViT-L achieved robust generalizability, our study did not evaluate the best-performing ViT-giant (ViT-g) architecture in DINOv2, a larger model with 1.1B parameters, which we expect to achieve better generalization performances in CXR interpretation, but it demands more pretraining data and computational resources.
@@
 6.3 Conclusion
-In summary, this work introduces CheXFound, a vision-centric foundation model pretrained via self-distillation on over one million unique CXRs.
-For downstream adaptations, we trained a GLoRI module on top of the frozen CheXFound, which combines disease-specific local features and global image features to improve the multilabel classification performance.
-CheXFound outperformed previous methods for classifying 40 disease findings on CXR-LT 24, demonstrating superior label efficiency on datasets with limited training labels, and strong generalization capabilities for opportunistic CXR interpretation on PLCO.
-The disease-specific local features extracted from CheXFound also carry strong interpretability as visuliazed by the associated attention maps.
-In our future work, we will continue to explore novel pretraining schemes to further improve the understanding of CXRs by these foundation models.
+In summary, this work introduces CheXFound, a vision-centric foundation model pretrained via self-distillation on over a large cohort of unique CXRs.
+For downstream adaptations, we trained a GLoRI head on top of the frozen CheXFound, which combined fine- and coarse-grained disease-specific local features and global image features to improve the multilabel classification performance.
+CheXFound outperformed previous methods for classifying 40 disease findings on CXR-LT 24.
+CheXFound demonstrated superior label efficiency on datasets with limited training labels and strong generalization capabilities to adapt to opportunistic CXR interpretation, malposition tube detection, and anatomical structure segmentation.
+The disease-specific local features extracted from CheXFound also manifested strong interpretability as characterized by the attention maps.
+In our future work, we will continue to explore novel pretraining schemes to further improve the understanding of CXRs by the foundation models.